### v1

# Fan distributions via Tverberg partitions and Gale duality

Abstract.
Equipartition theory, beginning with the classical ham sandwich theorem, seeks the fair division of finite point sets in $\mathbb{R}^{d}$ by full-dimensional regions determined by a prescribed geometric dissection of $\mathbb{R}^{d}$. Here we examine equidistributions of finite point sets in $\mathbb{R}^{d}$ by prescribed low dimensional subsets. Our main result states that if $r\geq 3$ is a prime power, then for any $m$-coloring of a sufficiently small point set $X$ in $\mathbb{R}^{d}$ there exists a certain type of $r$-fan in $\mathbb{R}^{d}$ – namely, the union of $r$ “half-flats” of codimension $r-2$ centered about a common $(r-1)$-codimensional affine subspace – which captures all the points of $X$ in such a way that each half-flat contains at most an $r$-th of the points from each color class. The number of points in $\mathbb{R}^{d}$ we require for this is essentially tight when $m\geq 2$, and for typical point sets the interiors of the half-flats must share a robust fraction of the points of $X$. We also give analogous equidistribution results for small point sets in $\mathbb{C}^{d}$ by regular $r$-fans which are centered about a complex hyperplane.
In the same way that a theorem of Dolnikov on hyperplane piercings generalizes the ham sandwich theorem, our equidistribution results extend to “piercing distributions” for families of subsets of $X$ satisfying particular intersection criteria. Our proof method continues recent work of Frick et al., where the Gale transform was shown to link Dolnikov-type hyperplane transversals and Radon-type partitions. Similarly, our results arise as the real and complex Gale duals of linear versions of Tverberg-type theorems. Finally, we extend our distribution results to multiple $r$-fans after establishing a multiple intersection version of a topological Tverberg-type theorem due to Sarkaria.
1. Introduction and Statement of Results
1.1. From Equipartitions to Equidistributions of Colored Point Sets
In recent years, equipartition problems have come to occupy a central place in geometric and topological combinatorics. Given a fixed integer $r\geq 2$ and a finite collection of finite point sets (or finite continuous measures) in $\mathbb{R}^{d}$, one seeks a prescribed geometric division of $\mathbb{R}^{d}$ into $r$ interior disjoint and typically convex regions, each of whose interiors contains no more than $1/r$ of the total number of points from each point set (respectively, of each total measure). The prototypical such result is the classical ham sandwich theorem [5, 31], which asserts that any $d$ finite points sets (or finite continuous measures) in $\mathbb{R}^{d}$ can be equipartitioned by a hyperplane, that is, by the two open half-spaces determined by the hyperplane. We refer the reader to the excellent recent survey [26] for more on equipartition theory.
Instead of equipartitions of finite point sets in $\mathbb{R}^{d}$ by full dimensional regions, here we seek their equidistribution by small-dimensional subsets of $\mathbb{R}^{d}$ arising from families of hyperplane with prescribed intersections. Specifically, suppose that $1\leq k\leq\min\{d-1,r-1\}$ and that $H_{1},\ldots,H_{r}$ are hyperplanes in $\mathbb{R}^{d}$ such that (1) the $r$-fold intersection of the hyperplanes is a flat of codimension $k+1$ and (2) any $k+1$ of the $r$ hyperplanes are independent (i.e., their normal vectors are linearly independent). Taking indices modulo $r$, for each $j\in[r]$ let $A_{j}$ be the affine space in $\mathbb{R}^{d}$ of codimension $k$ given by the $k$-fold intersection $A_{j}=\cap_{i=j+1}^{j+k}H_{i}$ of the hyperplanes and let $B_{j}=H_{j}^{+}\cap A_{j}$ be one of the (closed) half-flats of $A_{j}$ determined by a half-space $H_{j}^{+}$ of $H_{j}$. We call the union $F_{r}=\cup_{j=1}^{r}B_{j}$ of these half-flats an $r$-fan of codimenison $k$. The boundary of each half-flat is the common intersection $C=\cap_{i=1}^{r}H_{i}$ of the $r$ hyerplanes, which we call the center of the fan. In particular, an $r$-fan of codimension one is the union of $r$ half-hyperplanes centered about an affine space of codimension two. Equipartitions by the regions determined by such $r$-fans have been previously considered (see, e.g., [2, 3, 30]). We say that an $r$-fan of codimension one is regular if the dihedral angle between each of its successive half-hyperplanes is equal to $2\pi/r$.
Now suppose that $X$ is a finite set in $\mathbb{R}^{d}$. We say that an $r$-fan $F_{r}=\cup_{j=1}^{r}B_{j}$ in $\mathbb{R}^{d}$ distributes X if it contains $X$. If $X=X_{1}\sqcup\cdots\sqcup X_{m}$ is a partition of $X$ into $m$ non-empty subsets, which we call an $m$-coloring of $X$, we say that an $r$-fan equidistributes the coloring if it distributes $X$ and if the interior $\operatorname{\mathrm{Int}}(B_{j})=B_{j}\setminus C$ of each half-flat contains at most $1/r$ of the points from each $X_{i}$.
We have the following equidsitribution result for $r$-fans of codimension $r-2$. Recall that a set of points in $\mathbb{R}^{d}$ is affinely spanning if it is not contained in a hyperplane.
Theorem 1.1.
Let $d,m\geq 1$ be integers, let $r\geq 3$ be a prime power, and let $n\geq(r-1)(d+m+1)+1$. Then any $m$-coloring of a set of $n$ affinely spanning points in $\mathbb{R}^{n-d-1}$ can be equidistributed by an $r$-fan of codimension $r-2$.
Our proof of Theorem 1.1 shows that the interior of each half-flat of the equidistributing $r$-fan must contain at least one point from $X$, and, as we show in Section 5.1, for typical point sets the interiors of the half-flats must share at least $(r-1)(d+1)+1$ points of $X$. Similar remarks apply to Theorems 1.3, 1.4, and 1.6 below. As an example, letting $r=4$ and $d=m=1$ in Theorem 1.1 shows that for a typical set $X$ of $10$ points in $\mathbb{R}^{8}$ there is a $6$-dimensional $4$-fan which catches all $10$ points of $X$ so that either (a) $2$ points of $X$ lie in the fan’s $5$-dimensional center while the remaining $8$ points are equally dividing among the fan’s $4$ open half-flats, or (b) $3$ points lie in the center while the remaining $7$ are divided among its open half-flats so that each receives at most $2$ points.
Proposition 5.3 below will show that Theorem 1.1 is nearly sharp with respect to dimension for all fixed $r\geq 3$ and $m\geq 2$. Namely, for each integer $d\geq r-1$, let $n(r,m,d)$ be the maximum integer $n$ such that any $m$-coloring of any affinely spanning set of $n$ points in $\mathbb{R}^{d}$ can be equidistributed by an $r$-fan of codimension $r-2$. Once $d$ is sufficiently large, the lower bound on $n(r,m,d)$ given by Theorem 1.1 exceeds the exact value by at most two:
Corollary 1.2.
Let $r\geq 3$ be a prime power, let $m\geq 2$ be an integer, let $c=(r-1)(m+1)$, and let $d$ be an integer of the form $d=(r-2)s+t+c$ for some $s\geq 1$ and $t\in\{0,\ldots,r-2\}$. Then
$d+s\leq n(r,m,d)\leq d+s+2.$
When $r=3$, the $r$-fans of Theorem 1.1 have codimension one but need not be regular. If one considers point sets in $\mathbb{C}^{d}$, however, then for any prime power $r$ one can guarantee equidistributions by a complex regular $r$-fan in $\mathbb{C}^{d}$, that is a regular $r$-fan in $\mathbb{R}^{2d}$ of codimension one whose center is a complex hyperplane. For this, we require that the points complex affinely span $\mathbb{C}^{d}$, or in other words that they do not all lie on a complex hyperplane.
Theorem 1.3.
Let $d,m\geq 1$ be integers, let $r\geq 2$ be a prime power, and let $n\geq(r-1)(2d+m+1)+1$. Then any $m$-coloring of a set of $n$ complex affinely spanning points in $\mathbb{C}^{n-d-1}$ can be equidistributed by a complex regular $r$-fan in $\mathbb{C}^{n-d-1}$.
As with Theorem 1.1, the $r=2$ case of Theorem 1.3 is nearly sharp with respect to complex dimension (see Remark 5.4 below). As a regular $2$-fan in $\mathbb{R}^{d}$ is a hyperplane, the $r=2$ case of Theorem  1.3 is closely aligned with the ham sandwich theorem: for any $m$-coloring $X=X_{1}\sqcup\cdots\sqcup X_{m}$ of a set $X$ of $n$ complex affinely spanning points in $\mathbb{C}^{n-d-1}\cong\mathbb{R}^{2n-2d-2}$, there is a real hyperplane $H$ in $\mathbb{C}^{n-d-1}$ containing $X$ and a complex hyperplane $H_{\mathbb{C}}$ containing $H$ such that $H_{\mathbb{C}}$, when viewed as a real hyperplane in $H$, equipartitions each of the $X_{i}$.
1.2. Piercing Distributions
Recall that a family $\mathcal{F}$ of subsets of a finite set $X$ in $\mathbb{R}^{d}$ is said to be intersecting if any two members of $\mathcal{F}$ have non-empty intersection. The discrete case of a well-known theorem of Dolnikov [9] states that if $\mathcal{F}_{1},\ldots,\mathcal{F}_{d}$ are intersecting families of subsets of a finite set in $\mathbb{R}^{d}$, then there exists a hyperplane in $\mathbb{R}^{d}$ which intersects the convex hull of every member of each family. This result easily implies the ham sandwich theorem (see, e.g., [9, 12]). For prime powers $r\geq 3$, we give distribution results which are similar in spirit to Dolnikov’s theorem under the weaker intersection criterion that no family contains $r$ pairwise disjoint sets. Instead of a hyperplane which pierces the convex hull of every member of each family, we are guaranteed a distributing $r$-fan such every member of each family is itself pierced by two of the fan’s half-flats.
In what follows, it will be convenient to reformulate our intersection condition in terms of hypergraph colorings. Given a finite family $\mathcal{F}$ of non-empty subsets of a set $X$, its $r$-uniform Kneser hypergraph $\mathrm{KG}^{r}(\mathcal{F})$ is the $r$-uniform hypergraph whose vertices are the elements of $\mathcal{F}$ and whose hyperedges $\{F_{1},\ldots,F_{r}\}$ consist of $r$ pairwise-disjoint elements from $\mathcal{F}$. The chromatic number $\chi(\mathrm{KG}^{r}(\mathcal{F}))$ of $\mathrm{KG}^{r}(\mathcal{F})$ is the minimum number of colors needed to color all the vertices of $\mathrm{KG}^{r}(\mathcal{F})$ so that no hyperedge is monochromatic. Thus to say that $\chi(\mathrm{KG}^{r}(\mathcal{F}))\leq m$ means that there are $m$ non-empty subfamilies $\mathcal{F}_{1},\ldots,\mathcal{F}_{m}$ of $\mathcal{F}$ such that $\mathcal{F}=\cup_{j=1}^{r}\mathcal{F}_{i}$ and no $\mathcal{F}_{i}$ contains $r$ pairwise disjoint sets.
Theorem 1.4.
Let $d,m\geq 1$ be integers and let $r$ be a prime power.
  1. (a)
    Suppose that $r\geq 3$ and that $X$ is a set of $n\geq(r-1)(d+1)+1$ affinely spanning points in $\mathbb{R}^{n-d-1}$. If $\mathcal{F}$ is a family of non-empty subsets of $X$ with $\chi(\mathrm{KG}^{r}(\mathcal{F}))\leq m$, then there is an $r$-fan of codimension $r-2$ which distributes $X$ so that every $F\in\mathcal{F}$ is intersected by at least two closed half-flats of $F_{r}$.
  2. (b)
    Suppose that $r\geq 2$ and that $X$ is a set of $n\geq(r-1)(2d+m+1)+1$ complex affinely spanning points in $\mathbb{C}^{n-d-1}$. If $\mathcal{F}$ is a family of non-empty subsets of $X$ with $\chi(\mathrm{KG}^{r}(\mathcal{F}))\leq m$, then there is a complex regular $r$-fan $F_{r}$ in $\mathbb{C}^{n-d-1}$ which distributes $X$ so that every $F\in\mathcal{F}$ is intersected by at least two closed half-hyperplanes of $F_{r}$.
The $r=2$ case of Theorem 1.4(b) gives an analogue of Dolnikov’s theorem for complex instead of real hyperplanes: If $X$ is a set of $n$ complex affinely spanning points in $\mathbb{C}^{n-d-1}$ and $\mathcal{F}_{1},\ldots,\mathcal{F}_{m}$ are non-empty intersecting families of subsets of $X$, then there exists a real hyperplane $H$ in $\mathbb{C}^{n-d-1}$ which contains $X$ and a complex hyperplane $H_{\mathbb{C}}$ contained in $H$ which intersects the convex hull of every member of each $\mathcal{F}_{i}$.
1.3. Proof Scheme
We provide a brief preview of our proof method, which is a continuation of that used in  [12] to link Dolnikov-type transversal theorems with topological extensions of the classical Radon’s theorem [24]. Following this approach, Theorem 1.4, and in turn Theorems 1.1 and  1.3, is obtained as a consequence of a topological version of Tverberg’s celebrated generalization of Radon’s theorem.
Recall that Tverberg’s theorem [32] asserts that any set of $n\geq(r-1)(d+1)+1$ points in $\mathbb{R}^{d}$ can be partitioned into $r$ pairwise disjoint sets whose convex hulls have non-empty $r$-fold intersection. The $r=2$ case is Radon’s theorem. For prime powers, one has a topological extension of Tverberg theorem [22, 33] which asserts that any continuous map $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$ from the $(n-1)$-dimenisional simplex to $\mathbb{R}^{d}$ has an $r$-Tverberg tuple – that is, an $r$-tuple $(\sigma_{1},\ldots,\sigma_{r})$ of non-empty pairwise disjoint faces of $\Delta_{n-1}$ such that $f(\sigma_{1})\cap\cdots\cap f(\sigma_{r})\neq\emptyset$. When $r$ is a prime power, Tverberg’s theorem is recovered by letting $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$ be a linear map, that is, one which sends each convex combination of vertices of the simplex to the corresponding convex combination of the image of the vertices. In what follows, we identify the vertex set of $\Delta_{n-1}$ with $[n]=\{1,\ldots,n\}$ and we let $\operatorname{\mathrm{Vert}}(\sigma)\subseteq[n]$ denote the vertex set of any non-empty face $\sigma$ of $\Delta_{n-1}$.
As in  [12], we will apply the Gale transform, famous for its use in classification of polytopes with few vertices [10], to the linear cases of topological Tverberg-type theorems. In particular, Theorem 1.4 is in essence the Gale dual of the linear version of the following topological Tverberg-type theorem due to Sarkaria [27, 28].
Theorem 1.5.
Let $r\geq 2$ be a prime power, and let $d,m,n\geq 1$ be integers with $n\geq(r-1)(d+m+1)+1$. Suppose that $\mathcal{F}$ is a family of subsets of $[n]$ such that $\chi(\mathrm{KG}^{r}(\mathcal{F}))\leq m$. Then for any continuous map $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$, there exists an $r$-Tverberg-tuple $(\sigma_{1},\ldots,\sigma_{r})$ for $f$ such that no element of $\mathcal{F}$ is a subset of $\operatorname{\mathrm{Vert}}(\sigma_{j})$ for any $j\in[r]$.
As a further example of fan distributions obtainable from Tverberg-type results, we show that the Gale dual of the of the “optimal colored Tverberg theorem” gives rainbow distributions of a coloring of a finite point set, that is an $r$-fan which distributes the point set so that each of the fan’s open half-flats contains at most one point from each color class.
Theorem 1.6.
Let $d\geq 1$ be an integer.
  1. (a)
    Let $r\geq 4$ be an integer such that $r+1$ is prime and let $X$ be a set of $n\geq r(d+1)$ affinely spanning points in $\mathbb{R}^{n-d-1}$. If $X=X_{1}\sqcup\cdots\sqcup X_{d+1}$ is a $(d+1)$-coloring with $|X_{i}|\geq r$ for all $i\in[d+1]$, then there exists an $r$-fan of codimension $r-2$ which rainbow distributes the $(d+1)$-coloring.
  2. (b)
    Let $r\geq 2$ be an integer such that $r+1$ is prime and let $X$ be a set of $n\geq r(2d+1)-1$ complex affinely spanning points in $\mathbb{C}^{n-d-1}$. If $X=X_{1}\sqcup\cdots\sqcup X_{2d+1}$ is a $(2d+1)$-coloring of $X$ such that $|X_{i}|\geq r$ for all $i\in[2d+1]$, then there exists a complex regular $r$-fan in $\mathbb{C}^{n-d-1}$ which rainbow distributes the $(2d+1)$-coloring.
1.4. Distributions by Multiple Fans
Arguably the most fundamental problem in equipartition theory, dating back to Grünbaum [14] and Hadwiger [15], seeks the minimum dimension $d:=\Delta(m,k)$ such that any $m$ finite point sets in $\mathbb{R}^{d}$ can be equipartitioned by the $2^{k}$ regions determined by $k\geq 2$ independent hyperplanes. While a degrees of freedom argument  [25] gives the lower bound $\Delta(m,k)\geq\frac{m(2^{k}-1)}{k}$, very few tight upper bounds are currently known. For two hyperplanes in particular, sharp results exist only when either $m-1$, $m$, or $m+1$ is a power of two [7, 6, 19], in which case $\Delta(m,2)=\lfloor 3m/2\rfloor$. When $r$ is an odd prime, we give analogous extensions of our equidistribution results for two fans, where now given an $m$-coloring $X=X_{1}\sqcup\cdots\sqcup X_{m}$ of a finite point set $X$ in $\mathbb{R}^{d}$ we ask for two $r$-fans $F_{r}^{1}=\cup_{i=1}^{r}B^{1}_{i}$ and $F_{r}^{2}=\cup_{j=1}^{r}B^{2}_{j}$ in $\mathbb{R}^{d}$ whose intersection contains $X$ so that each intersection of the form $\operatorname{\mathrm{Int}}(B_{i}^{1})\cap\operatorname{\mathrm{Int}}(B_{j}^{2})$ contains at most $1/r^{2}$ of the points from each $X_{k}$. As with the calculations of the tight values of $\Delta(m,2)$, our results are ultimately derived from cohomological techniques which impose restrictions on the values of $m$.
Theorem 1.7.
Let $d\geq 1$ be an integer, let $r\geq 3$ be a prime number, and let $m\geq 1$ be an integer such that each of the coefficients in the $r$-ary expansion of $m(r-1)/2$ is even.
  1. (a)
    If $n\geq(r-1)(d+1)+\frac{m(r^{2}-1)}{2}+1$, then any $m$-coloring of a set $X$ of $n$ affinely spanning points in $\mathbb{R}^{n-d-1}$ can be equidistributed by two $r$-fans of codimension $r-2$.
  2. (b)
    If $n\geq(r-1)(2d+1)+\frac{m(r^{2}-1)}{2}+1$, then any $m$-coloring of a set of $X$ of $n$ affinely spanning points in $\mathbb{C}^{n-d-1}$ can be equidistributed by two complex regular $r$-fans in $\mathbb{C}^{n-d-1}$.
It is straightforward to show (see Remark 6.2) that the $r$-array expansion of $m(r-1)/2$ has all even coefficients provided $m$ is of the form $m=2a(r^{\ell_{1}}+\cdots+r^{\ell_{k}})$ where the $\ell_{i}\geq 0$ are distinct and $11\leq a\leq r-$ 1 is odd. In particular, one may let $m=2a(r^{\ell}-1)/(r-1)$ for any $\ell\geq 1$.
As was the case for a single fan, Theorem 1.7 is the special case of a Dolnikov-type piercing distribution result for two fans (see Theorem 8.1 below) which in turn is the Gale dual of a linear version of a multiple intersection version of Sarkaria’s theorem. In essence, this result guarantees two Tverberg pairs which are maximally distinct (see  [12] for the corresponding extension of Theorem 1.5 when $r=2$).
Theorem 1.8.
Let $d,m,r,n\geq 1$ be integers where $r\geq 3$ is a prime number, the $r$-ary coefficients of $m(r-1)/2$ are all even, and $n\geq(r-1)(d+1)+\frac{m(r^{2}-1)}{2}+1$. Suppose that $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$ is a continuous map and that $\mathcal{F}$ is a family of non-empty subsets of $[n]$ such that $\chi(\mathrm{KG}^{r^{2}}(\mathcal{F}))\leq m$. Then there exists two Tverberg-tuples $(\sigma_{1},\ldots,\sigma_{r})$ and $(\tau_{1},\ldots,\tau_{r})$ for $f$ such that no member of $\mathcal{F}$ is a subset of $\operatorname{\mathrm{Vert}}(\sigma_{i}\cap\tau_{j})$ for any $i,j\in[r]$.
Our proof of Theorem 1.8 follows essentially standard methods in topological combinatorics. In this case, the result reduces to a Borsuk-Ulam type statement for the group $\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}$ (Proposition 6.1) which we establish using a Chern class argument.
2. Real and Complex Gale Duality
In this section we discuss the properties of the real and complex Gale transform which we will need in translating between Tverberg partitions and fan distributions. These are quite standard in the real setting (see, e.g., [20]) and only require the minor adjustment of taking conjugates to hold in the complex setting. Nonetheless, for the sake of completeness we shall include their proofs here.
First, we remark that notions of affine dependence and affine span carry over from the real to the complex setting without issue. Thus a sequence of points $a_{1},\ldots,a_{n}$ in $\mathbb{C}^{d}$ is (complex) affinely dependent if there exists complex scalars $\lambda_{1},\ldots,\lambda_{n}$, not all zero, which simultaneously satisfy the conditions $\sum_{i=1}^{n}\lambda_{i}a_{i}=0$ and $\sum_{i=1}^{n}\lambda_{i}=0$. Likewise, the (complex) affine span of $a_{1},\ldots,a_{n}$ is defined to be the set of all affine combinations of the $a_{i}$, i.e. all the sums of the form $\sum_{i=1}^{n}\lambda_{i}a_{i}$ where $\sum_{i=1}^{n}\lambda_{i}=1$. These points affinely span $\mathbb{C}^{d}$ provided their affine span is all of $\mathbb{C}^{d}$.
Now let $\mathbb{K}$ denote the field of real or complex numbers and let $a_{1},\ldots,a_{n}$ be a sequence of $n$ points in $\mathbb{K}^{d}$ which ($\mathbb{K}$-) affinely span $\mathbb{K}^{d}$ (thus $n\geq d+1$). The Gale transform of this sequence is obtained as follows. First, let $A$ be the $(d+1)\times n$ matrix with coefficients in $\mathbb{K}$ whose column vectors are the $\widehat{a_{j}}=(a_{j},1)\in\mathbb{K}^{d+1}$. Because the $a_{j}$ affinely span $\mathbb{K}^{d}$, the $\widehat{a_{j}}$ linearly span $\mathbb{K}^{d+1}$. Thus $A$ has maximal rank and so its kernel has dimension $n-d-1$. Letting $b_{1},\ldots,b_{n-d-1}$ be a basis for $\ker A$, we let $B$ be the $(n-d-1)\times n$ matrix whose $i$-th row is $b_{i}$, and we let $\overline{B}$ denote its conjugate matrix. The Gale transform of $a_{1},\ldots,a_{n}$ is then defined to be the columns $g_{1},\ldots,g_{n}\in\mathbb{K}^{n-d-1}$ of $\overline{B}$. While the Gale transform is defined only up to a choice of basis for $\ker A$, the basic properties we shall need hold independently of such a choice, and moreover is not difficult to show that a different choice of basis corresponds to a change of the $g_{i}$ by a linear isomorphism of $\mathbb{K}^{n-d-1}$.
Proposition 2.1.
Let $g_{1},\ldots,g_{n}$ in $\mathbb{K}^{n-d-1}$ be the Gale transform of a sequence of $n$ affinely spanning points $a_{1},\ldots,a_{n}$ in $\mathbb{K}^{d}$. Then (1) the $g_{i}$ linearly span $\mathbb{K}^{n-d-1}$ and (2) $\sum_{i=1}^{n-d-1}g_{i}=0$. Conversely, any sequence $g_{1},\ldots,g_{n}$ in $\mathbb{K}^{n-d-1}$ satisfying (1) and (2) is the Gale dual of some sequence $a_{1},\ldots,a_{n}$ of affinely spanning points in $\mathbb{K}^{d}$.
Proof of Proposition 2.1.
For (1), note that if $b_{1},\ldots,b_{n-d-1}\in\mathbb{K}^{d+1}$ is a basis for $\ker A$, then the conjugate vectors $\overline{b}_{1},\ldots,\overline{b}_{n-d-1}$ are linearly independent in $\mathbb{K}^{d+1}$ as well. Thus $\overline{B}$ has maximal rank and so its columns linearly span $\mathbb{K}^{n-d-1}$. For (2), consider the conjugate $\bar{A}$ of $A$. As $Ab_{j}=0$ for all $j\in[n-d-1]$, $\bar{A}\,\overline{b_{j}}=0$ for all $j\in[n-d-1]$ as well. Since the last row of $\bar{A}$ is the all 1 vector, $\sum_{j=1}^{n}g_{j}=0$.
Now suppose that $g_{1},\ldots,g_{n}\in\mathbb{K}^{n-d-1}$ satisfies (1) and (2). Let $B$ be the $(n-d-1)\times n$ matrix whose columns are the $g_{i}$ and let $b_{1},\ldots,b_{n-d-1}$ be its rows. As the $g_{i}$ linearly span $\mathbb{K}^{n-d-1}$, $B$ has maximum rank and so $\dim\ker B=d+1$. Since $\sum_{i}g_{i}=0$, one can choose a basis $w_{1},\ldots,w_{d+1}$ for $\ker B$ with $w_{d+1}$ the all one vector. Now define $A$ to be the matrix whose rows are the $w_{i}$ and consider its conjugate $\bar{A}$. The columns of $A$ span $\mathbb{K}^{d+1}$ and therefore so do those of $\bar{A}$. As the last row of $\bar{A}$ is the all 1 vector, the columns of $\bar{A}$ are of the form $\widehat{a_{1}}=(a_{1},1),\ldots,\widehat{a_{n}}=(a_{n},1)$. Thus $a_{1},\ldots,a_{n}$ affinely span $\mathbb{K}^{d}$. By construction, we have $\ker\bar{A}=(\text{Row}\,A)^{\perp}=(\ker B)^{\perp}=\text{Row}\,\overline{B}.$ Thus $\overline{b}_{1},\ldots,\overline{b}_{n-d-1}$ is a basis for $\ker\bar{A}$, and so the columns $g_{1},\ldots,g_{n}$ of $\overline{\overline{B}}=B$ are the Gale transform of $a_{1},\ldots,a_{n}$. ∎
For our purposes, the essential property of the Gale transform is that it gives a correspondence between any affine dependency among the $a_{1},\ldots,a_{n}$ and an allocation of the $g_{1},\ldots,g_{n}$ with respect to a codimenison one linear subspace. More precisely, let $\langle u,w\rangle=\sum_{i=1}^{n}u_{i}\overline{w}_{i}$ denote the standard Hermitian inner product of two vectors $u$ and $w$ in $\mathbb{K}^{n}$. One then has the following.
Proposition 2.2.
Let $g_{1},\ldots,g_{n}\in\mathbb{K}^{n-d-1}$ be the Gale transform of a sequence $a_{1},\ldots,a_{n}$ of affinely spanning points in $\mathbb{K}^{d}$. If $\lambda=(\lambda_{1},\ldots,\lambda_{n})$ is a non-zero vector in $\mathbb{K}^{n}$ such that (1) $\sum_{i=1}^{n}\lambda_{i}a_{i}=0$ and (2) $\sum_{i=1}^{n}\lambda_{i}=0$, then there exists a non-zero vector $\alpha$ in $\mathbb{K}^{n-d-1}$ such that $\langle\alpha,g_{i}\rangle=\lambda_{i}$ for all $i\in[n-d-1]$. Conversely, given a non-zero vector $\alpha$ in $\mathbb{K}^{n-d-1}$, define $\lambda_{i}=\langle\alpha,g_{i}\rangle$ for all $i\in[n]$. Then $\lambda=(\lambda_{1},\ldots,\lambda_{n})\in\mathbb{K}^{n}$ is non-zero, $\sum_{i=1}^{n}\lambda_{i}a_{i}=0$, and $\sum_{i=1}^{n}\lambda_{i}=0$.
Proof of Proposition 2.2.
Let $A=\begin{bmatrix}\widehat{a_{1}}|&\cdots&|\widehat{a_{n}}\end{bmatrix}$ be the matrix associated to the $a_{i}$ in the construction of the Gale transform and let $b_{1},\ldots,b_{n-d-1}$ be a chosen basis for $\ker A$. First, suppose that $\lambda=(\lambda_{1},\ldots,\lambda_{n})\in\mathbb{K}^{n}\setminus\{0\}$ satisfies (1) and (2). Thus $\lambda\in\ker A$, and as $\lambda\neq 0$ there exists some $\alpha=(\alpha_{1},\ldots,\alpha_{n-d-1})\in\mathbb{K}^{n-d-1}\setminus\{0\}$ such that $\lambda=\sum_{j=1}^{n-d-1}\alpha_{j}b_{j}$. Taking the complex conjugate gives $\overline{\lambda_{i}}=\sum_{j=1}^{n-d-1}(\overline{b_{j}})_{i}\overline{\alpha_{j}}$ for all $i\in[n]$. But $(g_{i})_{j}=(\overline{b_{j}})_{i}$, so $\overline{\lambda_{i}}=\langle g_{i},\alpha\rangle$ and hence $\lambda_{i}=\langle\alpha,g_{i}\rangle$ as claimed. Conversely, suppose that $\alpha=(\alpha_{1},\ldots,\alpha_{n-d-1})\in\mathbb{K}^{n-d-1}\setminus\{0\}$. Defining $\lambda_{i}=\langle\alpha,g_{i}\rangle$ for all $i\in[n]$ and $\lambda=(\lambda_{1},\ldots,\lambda_{n})$, it follows easily from reversing the steps just given that $\lambda=\sum_{j=1}^{n-d-1}\alpha_{j}b_{j}$. Thus $\lambda\in\ker A$ and so (1) and (2) are satisfied. Finally, $\lambda\neq 0$ because the $b_{j}$ form a basis for $\ker A$.∎
3. Tverberg Partitions and Linear Fans
In this section, we describe an explicit connection between Tverberg $r$-tuples and linear $r$-fans, i.e, those whose center is a linear affine space. This connection will be used crucially in the proofs of Theorems 1.11.31.4, and 1.6 in Section 4.
Let $a_{1},\ldots,a_{n}$ be a sequence in $\mathbb{R}^{d}$ and let $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$ be the corresponding linear map determined by setting $f(i)=a_{i}$ for all $i\in[n]$. Supposing that $(\sigma_{1},\ldots,\sigma_{r})$ is a Tverberg $r$-tuple for $f$, we set $I_{j}=\operatorname{\mathrm{Vert}}(\sigma_{j})$ for all $j\in[r]$. We say that the Tverberg-tuple is proper if there exist non-negative real scalars $t_{1},\ldots,t_{n}$ such that
$\sum_{i\in I_{1}}t_{i}a_{i}=\cdots=\sum_{i\in I_{r}}t_{i}a_{i},$
where (1) $t_{i}>0$ for all $i\in\sqcup_{j=1}^{r}I_{j}$ and $\sum_{i\in I_{j}}t_{i}=1$ for all $j\in[r]$, and (2) $t_{i}=0$ for all $i\notin\sqcup_{j=1}^{r}I_{j}$.
3.1. Real Linear Fans
Given a sequence $x_{1},\ldots,x_{n}$ of $n$ points in $\mathbb{R}^{d}$ and $r$ non-empty pairwise disjoint subsets $I_{1},\ldots,I_{r}$ of $[n]$, we say an $r$-fan $F_{r}=\cup_{j=1}^{r}B_{j}$ in $\mathbb{R}^{d}$ with center $C$ distributes the $x_{i}$ according to the $I_{j}$ if
  1. (1)
    $x_{i}\in\operatorname{\mathrm{Int}}(B_{j})$ for all $i\in I_{j}$ and all $j\in[r]$, and
  2. (2)
    $x_{i}\in C$ for all $i\notin\sqcup_{j=1}^{r}I_{j}$.
Gale duality gives an equivalence between proper $r$-Tverberg partitions of sequences of points in $\mathbb{R}^{d}$ and distributions of sequences of points by linear $r$-fans of codimension $r-2$.
Lemma 3.1.
Let $d\geq 1$ and $r\geq 3$ be integers. Suppose that $a_{1},\ldots,a_{n}$ is a sequence of affinely spanning points in $\mathbb{R}^{d}$ with Gale transform $g_{1},\ldots,g_{n}$ in $\mathbb{R}^{n-d-1}$, and let $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$ be the affine linear map such that $f(i)=a_{i}$ for all $i\in[n]$.
  1. (a)
    Suppose that $(\sigma_{1},\ldots,\sigma_{r})$ is a proper $r$-Tverberg tuple for $f$ and let $I_{1},\ldots,I_{r}\subset[n]$ be the corresponding vertex sets of the $\sigma_{j}$. Then there exists a linear $r$-fan $F_{r}$ in $\mathbb{R}^{n-d-1}$ of codimension $r-2$ which distributes the $g_{i}$ according to the $I_{j}$.
  2. (b)
    Suppose that $I_{1},\ldots,I_{r}$ are non-empty pairwise disjoint subsets of $[n]$ and that $F_{r}$ is an $r$-fan in $\mathbb{R}^{n-d-1}$ of codimension $r-2$ which distributes the $g_{i}$ according to the $I_{j}$. Then there exists a proper $r$-Tverberg tuple $(\sigma_{1},\ldots,\sigma_{r})$ for $f$ such that $\operatorname{\mathrm{Vert}}(\sigma_{j})=I_{j}$ for all $j\in[r]$.
Proof of Lemma 3.1.
For part (a), suppose that $\sum_{i\in I_{1}}t_{i}a_{i}=\sum_{i\in I_{2}}t_{i}a_{i}=\cdots=\sum_{i\in I_{r}}t_{i}a_{i}$, where $\sum_{i\in I_{j}}t_{i}=1$ for all $j\in[r]$, $t_{i}>0$ for all $i\in\sqcup_{j=1}^{r}I_{j}$, and $t_{i}=0$ for all $i\notin\sqcup_{j=1}^{r}I_{r}$. For each $j\in[r]$, define $\lambda_{j}\in\mathbb{R}^{n}\setminus\{0\}$ by $(\lambda_{j})_{i}=t_{i}$ if $i\in I_{j}$, $(\lambda_{j})_{i}=-t_{i}$ if $i\in I_{j-1}$, and $(\lambda_{j})_{i}=0$ otherwise. Thus $\sum_{i=1}^{n}(\lambda_{j})_{i}=0$ and $\sum_{i=1}^{n}(\lambda_{j})_{i}a_{i}=0$. By Proposition 2.2, there is some non-zero vector $\alpha_{j}\in\mathbb{R}^{n-d-1}$ such that
$\langle\alpha_{j},g_{i}\rangle=(\lambda_{j})_{i}$ (3.1)
for all $i\in[n]$.
We claim that any $r-1$ of the $\alpha_{j}$ are linearly independent, but that $\alpha_{1},\ldots,\alpha_{r}$ are not. To see this, observe that since the $g_{i}$ linearly span $\mathbb{R}^{n-d-1}$, equation (3.1) implies that any linear relation $\sum_{j=1}^{n-d-1}\mu_{j}\alpha_{j}=0$ holds among the $\alpha_{j}$ if and only if the corresponding relation $\sum_{j=1}^{n-d-1}\mu_{j}\lambda_{j}=0$ holds among the $\lambda_{j}$. It is easily seen that any $r-1$ vectors among the $\lambda_{j}$ are linearly independent, so any $r-1$ of the $\alpha_{j}$ are as well. Similarly, $\sum_{j=1}^{r}\lambda_{j}=0$ implies that $\sum_{j=1}^{r}\alpha_{j}=0$. For each $j\in[r]$, let $H_{j}=\langle\alpha_{j}\rangle^{\perp}$ be the linear hyperplane corresponding to $\alpha_{j}$. Now let $A_{j}=\cap_{i\neq j,j-1}H_{i}$ and $B_{j}=H_{j}^{+}\cap A_{j}$, where $H_{j}^{+}=\{x\in\mathbb{R}^{n-d-1}\mid\langle\alpha_{j},x\rangle\geq 0\}$. Thus $H_{j}\cap A_{j}=\cap_{i=1}^{r}H_{i}$ for all $j\in[r]$ and $F_{r}=\cup_{j=1}^{r}B_{j}$ is a linear $r$-fan of codimension $r-2$. The distribution claim about the $g_{i}$ also follows from equation (3.1). For if $j\in[r]$ and $i\in I_{j}$, we have $\langle\alpha_{j},g_{i}\rangle=(\lambda_{j})_{i}=t_{i}>0$, while $\langle\alpha_{k},g_{i}\rangle=(\lambda_{k})_{i}=0$ if $k\neq j-1,j$. Thus $g_{i}\in\operatorname{\mathrm{Int}}(B_{j})=\operatorname{\mathrm{Int}}(H_{j}^{+})\cap A_{j}$. On the other hand, $\langle g_{i},\alpha_{j}\rangle=0$ for all $j\in[r]$ if $i\notin\sqcup_{j=1}^{r}I_{j}$ and therefore $g_{i}\in C$.
For part (b), suppose that $F_{r}$ a linear $r$-fan in $\mathbb{R}^{n-d-1}$ of codimenison $r-2$ determined by $H_{1}=\langle\alpha_{1}\rangle^{\perp},\ldots,H_{r}=\langle\alpha_{r}\rangle^{\perp}$ in $\mathbb{R}^{n-d-1}$. By assumption, $\alpha_{1},\ldots,\alpha_{r}$ are linearly dependent, so $\sum_{j=1}^{r}\mu_{j}\alpha_{j}=0$ for some scalars $\mu_{1},\ldots,\mu_{r}$, not all of which are zero. Since any $r-1$ of the $\alpha_{j}$ are linearly independent, $\mu_{j}\neq 0$ for all $j\in[r]$. Thus by scaling both $\alpha_{j}$ and $\beta_{j}$ by $\mu_{j}$ we may assume that $\sum_{j=1}^{r}\alpha_{j}=0$. Without loss of generality, we may let $H^{+}_{j}=\{x\in\mathbb{R}^{n-d-1}\mid\langle\alpha_{j},x\rangle\geq 0\}$ for all $j\in[r]$.
For each $j\in[r]$, define $\lambda_{j}=(\lambda_{j,1},\ldots,\lambda_{j,n})\in\mathbb{R}^{n}$ by $\lambda_{j,i}=\langle\alpha_{j},g_{i}\rangle$ for all $i\in[n]$. As $g_{i}\in\operatorname{\mathrm{Int}}(B_{j})$ for all $i\in I_{j}$, $\lambda_{j}$ is non-zero and so by Proposition 2.2 we have $\sum_{i=1}^{n}\lambda_{j,i}=0$ and $\sum_{i=1}^{n}\lambda_{j,i}a_{i}=0$. Now let $j\in[r]$ and suppose that $i\in I_{j}$. Since $g_{i}\in\operatorname{\mathrm{Int}}(B_{j})$, we have $\lambda_{j,i}>0$ and $\lambda_{k,i}=0$ for all $k\neq j-1,j$. As $\sum_{j=1}^{r}\alpha_{j}=0$, we have $\langle\sum_{j=1}^{r}\alpha_{j},g_{i}\rangle=\langle\alpha_{j-1},g_{i}\rangle+\langle\alpha_{j},g_{i}\rangle=0$. Thus $\lambda_{j-1,i}=-\lambda_{j,i}$ for all $i\in I_{j}$. Again fix $j\in[r]$ and consider the relation $\sum_{i=1}^{n}\lambda_{j,i}a_{i}=0$. If $i\notin I_{j}\sqcup I_{j+1}$, then $g_{i}\notin\operatorname{\mathrm{Int}}(B_{j})\sqcup\operatorname{\mathrm{Int}}(B_{j+1})$. In particular, $g_{i}\in\cap_{\ell\neq k,k-1}H_{\ell}$ for some $k\neq j,j+1$ and so $\langle\alpha_{j},g_{i}\rangle=0$ for all $i\notin I_{j}\sqcup I_{j+1}$. Thus $\sum_{i\in I_{j}}\lambda_{j,i}a_{i}+\sum_{i\in I_{j+1}}\lambda_{j,i}a_{i}=0$ and so $\sum_{i\in I_{j}}\lambda_{j,i}a_{i}=-\sum_{i\in I_{j+1}}\lambda_{j,i}a_{i}=\sum_{i\in I_{j+1}}\lambda_{j+1,i}a_{i}$ for any $j\in[r]$. We therefore have
$\sum_{i\in I_{1}}\lambda_{i}a_{i}=\cdots=\sum_{i\in I_{r}}\lambda_{i}a_{i}.$ (3.2)
For each $j\in[r]$ we let $s_{j}=\sum_{i\in I_{j}}\lambda_{j,i}$. As $\langle\alpha_{j},g_{i}\rangle>0$ for all $i\in I_{j}$, we have $s_{j}>0$. Letting $t_{i}=\lambda_{j,i}/s_{j}$ for each $i\in I_{j}$, we have that $t_{i}>0$ for all $i\in I_{j}$ and that $\sum_{i\in I_{j}}t_{i}=1$. Fixing $j\in[r]$ and considering the sum $\sum_{i=1}^{n}\lambda_{j,i}=0$, the identical reasoning to that used in deriving equation (3.2) shows that $\sum_{i\in I_{j}}\lambda_{j,i}=\sum_{i\in I_{j+1}}\lambda_{j+1,i}$ for all $j\in[r]$ and so that $s_{1}=\cdots=s_{r}$. Dividing equation (3.2) by $s_{1}$, we now have $\sum_{i\in I_{1}}t_{i}a_{i}=\cdots=\sum_{i\in I_{r}}t_{i}a_{i}$. This gives a proper Tverberg tuple for the linear map $f$. ∎
3.2. Complex Linear Fans
Before giving a complex analogue of Lemma 3.1, we first give a description of complex regular $r$-fans in $\mathbb{C}^{d}$ in terms of the standard Hermitian inner product. Any complex hyperplane in $\mathbb{C}^{d}$ is of the form $H_{\mathbb{C}}:=\{z\in\mathbb{C}^{d}\mid\langle\alpha,z\rangle=\beta\}$ for some $(\alpha,\beta)\in\mathbb{C}^{n}\setminus\{0\}\times\mathbb{C}$. Let $\omega_{r}=\exp(2\pi i/r)$ be the standard $r$-th root of unity. If $F_{r}=\cup_{j=1}^{r}B_{j}$ is a complex regular $r$-fan centered about $H_{\mathbb{C}}$, then there exists some unit complex number $\lambda$ such that $B_{j}=\{z\in\mathbb{C}^{d}\mid\langle\alpha,z\rangle=\beta+t\lambda\omega_{r}^{j}\,\,\text{for some}\,\,t>0\}$ for all $j\in[r]$. Moreover, one may take $\lambda=1$ after scaling $(\alpha,\beta)$ by $\lambda^{-1}$.
Lemma 3.2.
Let $d\geq 1$ and $r\geq 2$ be integers. Suppose that $a_{1},\ldots,a_{n}$ are complex affinely spanning points in $\mathbb{C}^{d}$ whose Gale transform is $g_{1},\ldots,g_{n}\in\mathbb{C}^{n-d-1}$, and let $f\colon\Delta_{n-1}\rightarrow\mathbb{C}^{d}\cong\mathbb{R}^{2d}$ be the (real) linear map given by setting $f(i)=a_{i}$ for all $i\in[n]$.Suppose that $(\sigma_{1},\ldots,\sigma_{r})$ is a proper $r$-Tverberg tuple for $f$ and let $I_{1},\ldots,I_{r}\subset[n]$ be the corresponding vertex sets of the $\sigma_{j}$. Then there exists a linear complex regular $r$-fan $F_{r}$ in $\mathbb{C}^{n-d-1}$ which distributes the $g_{i}$ according to the $I_{j}$.
Proof of Lemma 3.2.
As before, suppose that $I_{1},\ldots,I_{r}\subset[n]$ are pairwise disjoint non-empty sets such that $\sum_{i\in I_{1}}t_{i}a_{i}=\cdots=\sum_{i\in I_{r}}t_{i}a_{i}$ where $t_{i}>0$ for all $i\in\sqcup_{j=1}^{r}I_{j}$ and $\sum_{i\in I_{j}}t_{i}=1$ for all $j\in[r]$. To construct a complex affine dependence among the $a_{i}$, define $\lambda=(\lambda_{1},\ldots,\lambda_{n})\in\mathbb{C}^{n}\setminus\{0\}$ as follows: for each $j\in[r]$ and each $i\in I_{j}$, let $\lambda_{i}=t_{i}\omega_{r}^{j}$, and let $\lambda_{i}=0$ for all $i\notin\sqcup_{j=1}I_{j}$. As $\sum_{i\in I_{j}}t_{j}=1$ for all $j\in[r]$, we have $\sum_{i=1}^{n}\lambda_{i}=\sum_{j\in[r]}\sum_{i\in I_{j}}t_{i}\omega_{r}^{j}=\sum_{j\in[r]}\omega_{j}^{r}=0$. Likewise, we have $\sum_{i\in I_{j}}t_{i}a_{i}=c$ is constant over $j\in[r]$ and so $\sum_{i\in[n]}\lambda_{i}a_{i}=\sum_{j\in[r]}\sum_{i\in I_{j}}t_{i}\omega_{r}^{j}a_{i}=c\sum_{j\in[r]}\omega_{r}^{j}=0$ as well. By Proposition 2.2, there is some $\alpha\neq 0$ such that $\langle\alpha,g_{i}\rangle=\lambda_{i}$ for all $i\in[n]$. Let $H_{\mathbb{C}}=\langle\alpha\rangle^{\perp}$ be the resulting linear complex hyperplane in $\mathbb{C}^{n-d-1}$, and for each $j\in[r]$ let $B_{j}=\{z\in\mathbb{C}^{n-d-1}\mid\langle\alpha,z\rangle=t\omega_{r}^{j}\,\,\text{for some}\,\,t\geq 0\}$. Thus $F_{r}=\cup_{j=1}^{r}B_{j}$ is a complex regular $r$-fan $F_{r}$ centered about $H_{\mathbb{C}}$. By construction, $g_{i}\in\operatorname{\mathrm{Int}}{B_{j}}$ for all $i\in I_{j}$ and any $j\in[r]$, while $g_{i}\in H_{\mathbb{C}}$ if $i\notin\sqcup_{j\in[r]}I_{j}$. ∎
Remark 3.3.
We observe that distributions by complex linear $r$-fans are equivalent to proper Tverberg partitions when $r=2$. To see this, suppose that $H=B_{1}\cup B_{2}$ is a real hyperplane in $\mathbb{C}^{n-d-1}$ centered about a complex linear hyperplane $H_{\mathbb{C}}=\langle\alpha\rangle^{\perp}$ and that there exist two disjoint non-empty subsets $I_{1},I_{2}$ of $[n]$ such that $g_{i}\in\operatorname{\mathrm{Int}}{B_{j}}$ for all $i\in I_{j}$ and all $j\in\{1,2\}$, while $g_{i}\in H_{\mathbb{C}}$ otherwise. Thus $\langle\alpha,g_{i}\rangle=t_{i}(-1)^{j}$ for all $i\in I_{j}$ and all $j\in\{1,2\}$, where $t_{i}>0$ for all $i\in I_{1}\sqcup I_{2}$, while $\langle\alpha,g_{i}\rangle=0$ if $i\notin I_{1}\sqcup I_{2}$. By Proposition 2.2, we have $\sum_{i\in I_{1}}t_{i}x_{i}=\sum_{i\in I_{2}}t_{i}x_{i}$ and $\sum_{j\in I_{1}}t_{i}=\sum_{j\in I_{2}}t_{i}$. Letting $t_{i}^{\prime}=\frac{t_{i}}{\sum_{j\in I_{1}}t_{i}}$ for all $i\in I_{1}\sqcup I_{2}$, we have that $t_{i}^{\prime}>0$ for all $i\in I_{1}\sqcup I_{2}$, $\sum_{i\in I_{1}}t_{i}^{\prime}=\sum_{i\in I_{2}}t_{i}^{\prime}$, and $\sum_{i\in I_{1}}t_{i}^{\prime}x_{i}=\sum_{i\in I_{2}}t_{i}^{\prime}x_{i}$.
4. Distributions by a Single Fan
We now prove our distribution results for a single fan. Each of these follows by applying Gale duality to the linear versions of associated topological Tverberg-type theorem. We defer the proof of the near optimality of Theorem of the real case of Theorem 1.4) Section 5.2. Analogous remarks apply to the complex case of Theorem 1.3 when $r=2$.
4.1. Proof of Theorems 1.11.3, and  1.4
We begin by proving our equdistribution results for a single fan, which are a special case of Theorem 1.4.
Proof of Theorems 1.1 and  1.3.
Let $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$. We let $n\geq(r-1)(d+m+1)$ when $\mathbb{K}=\mathbb{R}$ and $n\geq(r-1)(2d+m+1)+1$ when $\mathbb{K}=\mathbb{C}$. Suppose that $X=X_{1}\sqcup\cdots\sqcup X_{m}$ is an $m$-coloring of a set of $n$ affinely spanning points in $\mathbb{K}^{n-d-1}$. Corresponding to each $X_{i}$, define a corresponding family of non-empty subsets of $X$ by $\mathcal{F}_{i}=\{F\subset X\mid|F\cap X_{i}|>|X_{i}|/r\}$ and let $\mathcal{F}=\cup_{i=1}^{m}\mathcal{F}_{i}$. As no $\mathcal{F}_{i}$ can contain $r$ pairwise disjoint members by the pigeon-hole principle, $\chi(\mathrm{KG}^{r}(\mathcal{F}))\leq m$. Theorem 1.4(a) guarantees an $r$-fan $F_{r}=\cup_{j=1}^{r}B_{j}$ in $\mathbb{K}^{n-d-1}$ (an $r$-fan of codimension $r-2$ when $\mathbb{K}=\mathbb{R}$, and a complex regular fan when $\mathbb{K}=\mathbb{C})$ such that no $\operatorname{\mathrm{Int}}(B_{j})$ contains any element of $\mathcal{F}$. It follows that $|\operatorname{\mathrm{Int}}(B_{j})\cap X_{i}|\leq|X_{i}|/m$ for every $j\in[r]$ and each $i\in[m]$, so that $F_{r}$ is the desired $r$-fan. Indeed, were $|\operatorname{\mathrm{Int}}(B_{j})\cap X_{i}|>|X_{i}|/m$ for some $j\in[r]$ and $i\in[m]$ we would have $\operatorname{\mathrm{Int}}(B_{j})\cap X\in\mathcal{F}_{i}$, and therefore that $\operatorname{\mathrm{Int}}(B_{j})\cap X$ is not a subset of $\operatorname{\mathrm{Int}}(B_{j})$. ∎
We now prove Theorem 1.4.
Proof of Theorem 1.4.
Again let $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$ and let $X=\{x_{1},\ldots,x_{n}\}$ be a set of $n$ affinely spanning points in $\mathbb{K}^{n-d-1}$, where $n\geq(r-1)(d+m+1)+1$ when $\mathbb{K}=\mathbb{R}$ and $n\geq(r-1)(2d+m+1)+1$ if $\mathbb{K}=\mathbb{C}$. We are given that $\mathcal{F}$ is a family of non-empty subsets of $X$ satisfying $\chi(\mathrm{KG}^{r}(\mathcal{F}))\leq m$. We now lift $X$ to the $\mathbb{K}$-hyperplane $\mathbb{K}^{n-d-1}\times\{1\}$ of height one in $\mathbb{K}^{n-d}$. The sequence $g_{1}=(x_{1},1),\ldots,g_{n}=(x_{n},1)$ in $\mathbb{K}^{n-d}$ is linearly spanning, and including $g_{n+1}=-\sum_{i=1}^{n}g_{i}$ gives a sequence of $n+1$ points in $\mathbb{K}^{n-d}$ which both linearly span and which sum to zero. By Proposition 2.1, the sequence $g_{1},\ldots,g_{n+1}$ is the Gale transform of $n+1$ points $a_{1},\ldots,a_{n+1}$ in $\mathbb{K}^{d}$. We let $f\colon\Delta_{n}\rightarrow\mathbb{K}^{d}$ be the (real) linear map determined by sending each $i\in[n+1]$ to $a_{i}$ (here we identify $\mathbb{C}^{d}$ with $\mathbb{R}^{2d}$).
Corresponding to $\mathcal{F}$, let $\mathcal{F}^{\prime}$ be the family of all subsets $\{i_{1},\ldots,i_{k}\}$ of $[n]$ such that $\{x_{i_{1}},\ldots,x_{i_{k}}\}$ is in $\mathcal{F}$. As $\chi(\mathrm{KG}^{r}(\mathcal{F}))\leq m$, $\chi(\mathrm{KG}^{r}(\mathcal{F}^{\prime}))\leq m$ as well. Given the value of $n$, we may apply Theorem 1.5 to the restriction $f\mid_{\Delta_{n-1}}\colon\Delta_{n-1}\rightarrow\mathbb{K}^{d}$ of $f$ to the simplex $\Delta_{n-1}$ whose vertex set is $[n]$ Thus there is a proper Tverberg pair $(\sigma_{1},\ldots,\sigma_{r})$ for $f|_{\Delta_{n-1}}$ such that each $I_{j}=\operatorname{\mathrm{Vert}}(\sigma_{j})\subset[n]$ does not contain any element of $\mathcal{F}^{\prime}$. As $(\sigma_{1},\ldots,\sigma_{r})$ is also a proper $r$-Tverberg partition for the original map $f$, by Lemma 3.1 (respectively, Lemma 3.2) there is a linear $r$-fan $F_{r}$ in $\mathbb{K}^{n-d}$ (of codimension $r-2$ when $r=2$ and complex linear when $\mathbb{K}=\mathbb{C}$) which distributes $g_{1},\ldots,g_{n+1}$ according to the $I_{j}$. Thus $g_{i}\in\operatorname{\mathrm{Int}}(B_{j})$ for all $i\in I_{j}$ and all $j\in[r]$, while $g_{i}\in C=\cap_{j=1}^{r}H_{j}$ for all $i\notin\sqcup_{j=1}^{r}I_{j}$. As $n+1$ does not lie in any of the $I_{j}$, we have that $g_{n+1}$ lies on $C$.
Now let $\mathbb{K}=\mathbb{R}$. For each $j\in[r]$ we have $H_{j}=\langle\alpha_{j}\rangle^{\perp}$ where $\alpha_{j}=(\beta_{j},\gamma_{j})\in\mathbb{R}^{n-d-1}\times\mathbb{R}$ is non-zero. The $r$-fan is of the form $F_{r}=\cup_{j=1}^{r}B_{j}$, where each $B_{j}$ is of the form $B_{j}=H_{j}^{+}\cap A_{j}$ with $A_{j}=\cap_{i\neq j,j-1}H_{j}$ and $H_{j}^{+}$ being some open half-space of $H_{j}^{+}$. For each $j\in[r]$, let $H^{\prime}_{j}=H_{j}\cap(\mathbb{R}^{n-d-1}\times\{1\})$ be the intersection of $H_{j}$ with the plane of height one in $\mathbb{R}^{n-d}$. Thus $H^{\prime}_{j}=\{(x,1)\in\mathbb{R}^{n-d-1}\times\{1\}\mid\langle\beta_{j},x\rangle=-\gamma_{j}\}$. Likewise, let $B^{\prime}_{j}=B_{j}\cap(\mathbb{R}^{n-d-1}\times\{1\})$ for all $j\in[r]$. Letting $H_{j}^{\prime\prime}$ and $B_{j}^{\prime\prime}$ be the corresponding projections of $H_{j}^{\prime}$ and $B_{j}^{\prime}$ onto $\mathbb{R}^{n-d-1}$, we will show that $F^{\prime\prime}_{r}=\cup_{j=1}^{r}B_{j}^{\prime\prime}$ is an $r$-fan of codimension $r-2$ in $\mathbb{R}^{n-d-1}$.
To that end, we first show that any $r-1$ of $\beta_{1},\ldots,\beta_{r}$ are linearly dependent in $\mathbb{R}^{n-d-1}$. In particular, each $\beta_{j}$ is non-zero, so each $H^{\prime\prime}_{j}$ is a hyperplane in $\mathbb{R}^{n-d-1}$. Without loss of generality, consider $\beta_{1},\ldots,\beta_{r-1}$ and suppose that $\sum_{j=1}^{r-1}\mu_{j}\beta_{j}=0$. We have $\langle\sum_{j=1}^{r-1}\mu_{j}\alpha_{j},g_{n+1}\rangle=0$ because $g_{n+1}\in H_{j}$ for all $j\in[r]$. As $g_{n+1}=(-x,-n)$ where $x=\sum_{i=1}^{n}x_{i}\in\mathbb{R}^{n-d-1}$, we have $\langle\sum_{j-1}^{r}\mu_{j}\alpha_{j},g_{n+1}\rangle=-n\sum_{j=1}^{r-1}\mu_{j}\gamma_{j}$. Thus $\sum_{j=1}^{r-1}\mu_{j}\gamma_{j}=0$, and since $\sum_{j=1}^{r-1}\mu_{j}\beta_{j}=0$ by assumption we have that $\sum_{j=1}^{r-1}\mu_{j}\alpha_{j}=0$ as well. However, any $r-1$ of the $\alpha_{j}$ are linearly independent in $\mathbb{R}^{n-d}$, so each $\mu_{j}$ is zero and $\beta_{1},\ldots,\beta_{r-1}$ are linearly independent in $\mathbb{R}^{n-d-1}$. On the other hand, $\beta_{1},\ldots,\beta_{r}$ are linearly dependent because $\alpha_{1},\ldots,\alpha_{r}$ are. Finally, we show that $\cap_{j=1}^{r}H^{\prime}_{j}\neq\emptyset$ and so that $H_{j}^{\prime\prime}\neq\emptyset$ as well. But $g_{n+1}=(-x,-n)\in\cap_{j=1}^{r}H_{j}$, so $(\frac{-x}{n},1)\in\cap_{j=1}^{r}H_{j}^{\prime}\neq\emptyset$.
In the complex setting, we have a complex regular $r$-fan $F_{r}=\cup_{j=1}^{r}B_{j}$ in $\mathbb{C}^{n-d}$ centered about a linear complex hyperplane $H_{\mathbb{C}}$. As above, we let $H_{\mathbb{C}}^{\prime}=H_{\mathbb{C}}\cap(\mathbb{C}^{n-d-1}\times\{1\})$, $B_{j}^{\prime}=B_{j}\cap(\mathbb{C}^{n-d-1}\times\{1\})$ for each $j\in[r]$, and we let $H^{\prime\prime}_{\mathbb{C}}$ and $B^{\prime\prime}_{j}$ be their corresponding projections onto $\mathbb{C}^{n-d-1}$. We claim that $F_{r}^{\prime\prime}=\cup_{j=1}^{r}B^{\prime\prime}_{j}$ is a regular $r$-fan in $\mathbb{C}^{n-d-1}$ centered about the complex hyerplane $H^{\prime\prime}_{\mathbb{C}}$. To verify this, we have $H_{\mathbb{C}}=\langle\alpha\rangle^{\perp}$, where $\alpha=(\beta,\gamma)\in\mathbb{C}^{n-d-1}\times\mathbb{C}$ is non-zero. Thus $H^{\prime\prime}_{\mathbb{C}}=\{z\in\mathbb{C}^{n-d-1}\mid\langle\beta,z\rangle=-\gamma\}$ and $B^{\prime}_{j}=\{(z\in\mathbb{C}^{n-d-1}\mid\langle\beta,z\rangle+\gamma=t\omega_{r}^{j}\,\,\text{for some}\,\,t\geq 0\}$. Thus $H_{\mathbb{C}}^{\prime\prime}$ and $F_{r}^{\prime\prime}$ are as claimed provided $\beta\neq 0$. But $g_{n+1}=(-x,-n)\in H_{\mathbb{C}}$ where $x=\sum_{i=1}^{n}g_{i}$. Thus $\langle(x,-n),(\beta,\gamma)\rangle=0$ and so $\langle\beta,x\rangle=n\gamma$. But if $\beta=0$, then $\gamma=0$ and therefore $\alpha=0$ as well. Thus $F^{\prime\prime}_{r}$ is indeed a regular $r$-fan centered about the complex hyperplane $H_{\mathbb{C}}^{\prime\prime}$.
To complete the proof, we again let $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$ and show that the interior $\operatorname{\mathrm{Int}}(B^{\prime\prime}_{j})$ of any $B^{\prime\prime}_{j}$ does not contain any element of $\mathcal{F}$. So let $F\in\mathcal{F}$ and let $F^{\prime}$ be the corresponding element of $\mathcal{F}^{\prime}$. We have that $|F^{\prime}|=|F|$ and $|F^{\prime}\cap I_{j}|=|F\cap\operatorname{\mathrm{Int}}(B^{\prime\prime}_{j})|$ for all $j\in[r]$. Since $F^{\prime}$ is not contained in any of the $I_{j}$, we conclude that $F$ is not contained in any of the $\operatorname{\mathrm{Int}}(B^{\prime\prime}_{j})$. ∎
4.2. Proof of Theorem 1.6
We conclude this section by proving our colorful distribution result. This follows the same procedure as before, except we substitue Sarkaria’s theorem with the Optimal Colored Tverberg theorem of  [8] below.
Theorem 4.1.
Let $d\geq 1$ be an integer, let $r\geq 2$ be an integer such that $r+1$ is prime, and let $n\geq r(d+1)$. If $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$ is a continuous map and $C_{1}\sqcup\cdots\sqcup C_{d+1}$ is a $(d+1)$-coloring of the vertex set of $\Delta_{n}$ such that each $|C_{i}|\geq r$ for all $i\in[d+1]$, then there exists a Tverberg $r$-tuple $(\sigma_{1},\ldots,\sigma_{r})$ for $f$ such that $|\sigma_{j}\cap C_{i}|\leq 1$ for all $j\in[r]$ and all $i\in[d+1]$.
Proof of Theorem 1.6.
Let $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$. We let $n\geq r(d+1)$ if $\mathbb{K}=\mathbb{R}$ and $n\geq r(2d+1)$ if $\mathbb{K}=\mathbb{C}$ and we suppose that $X=\{x_{1},\ldots,x_{n}\}$ is a set of $n$ affinely spanning points in $\mathbb{K}^{n-d-1}$. In the real case we have that $X=X_{1}\sqcup\cdots\sqcup X_{d+1}$ is a $(d+1)$-coloring with $|X_{k}|\geq r$ for all $k\in[d+1]$, and likewise in the complex case that with $d+1$ replaced by $12d+$ 1. As in the proof of Theorem 1.4, lift $X$ to the hyperplane $\mathbb{K}^{n-d-1}\times\{1\}$ of height $1$ in $\mathbb{K}^{n-d}$. As there, the points $g_{1}=(x_{1},1)\,\ldots,g_{n}=(x_{n},1),g_{n+1}=-\sum_{i=1}^{n}g_{i}$ are the Gale transform of some sequence $a_{1},\ldots,a_{n+1}$ of affinely spanning points in $\mathbb{K}^{d}$. As before, we let $f\colon\Delta_{n}\rightarrow\mathbb{K}^{d}$ be the (real) affine linear map defined by $f(i)=a_{i}$ for all $i\in[n+1]$. If $\mathbb{K}=\mathbb{R}$, we define $C_{k}=\{i\in[n]\mid x_{i}\in X_{k}\}$ for each $k\in[d+1]$. Thus $C_{1}\sqcup$ is a $(d+1)$-coloring of the vertex set of $\Delta_{n-1}$ into $d+1$ color classes of size at least $r$ each. We proceed analogously in the complex setting. Applying Theorem 4.1 to the restriction of $f$ to $\Delta_{n-1}$, we have a proper $r$-Tverberg tuple $(\sigma_{1},\ldots,\sigma_{r})$ for $f\mid_{\Delta_{n-1}}$ such that the vertex set $I_{j}\subset[n]$ of each $\sigma_{j}$ contains at most one element of each color class.
In the real case, let $H_{j}$, $B_{j}$, $H^{\prime\prime}_{j}$, and $B^{\prime\prime}_{j}$ be as in the proof of Theorem 1.4(a). The argument given there shows that $F^{\prime\prime}_{r}=\cup_{j=1}^{r}B^{\prime\prime}_{j}$ is an $r$-fan in $\mathbb{R}^{n-d-1}$ of codimension $r-2$ such that $x_{i}$ lies in $\operatorname{\mathrm{Int}}(B^{\prime\prime}_{j})$ for all $i\in I_{j}$ and all $j\in[r]$, while $x_{i}$ lies in the center of the fan for all $i\notin\sqcup_{j=1}^{r}I_{j}$. Likewise, let $F_{r}$ and $F_{r}^{\prime\prime}$ be as in the proof of Theorem 1.4(b). Thus $F_{r}^{\prime\prime}$ is a complex regular $r$-fan in $\mathbb{C}^{n-d-1}$ which also distributes $X$ according to the $I_{j}$. To complete the proof, we only need to show that $|\operatorname{\mathrm{Int}}(B_{j}^{\prime\prime})\cap X_{k}|\leq 1$ for all $j$ and $k$. But $|\operatorname{\mathrm{Int}}(B_{j}^{\prime\prime})\cap X_{k}|=|I_{j}\cap C_{k}|\leq 1$ for all $j$ and $k$, so the proof is complete.∎
5. Optimality of Fan Distributions and Robustness for Typical Point sets
In this section we show that there is a dense open class of affinely spanning point sets for which the distributing $r$-fans of Theorems 1.4 and 1.6 contain a robust fraction of the point set. In addition, we prove various cases of the near optimality of Theorem 1.4 with respect to dimension.
5.1. Typical Point Sets
Let $a_{1},\ldots,a_{n}$ be a sequence of $n$ affinely spanning points in $\mathbb{K}^{d}$ and let $a_{n+1}$ by the average of the $a_{i}$. Letting $A$ be the matrix of the Gale transform associated to $a_{1},\ldots,a_{n+1}$, one may choose the vector $(1,\ldots,1,-n)$ to be in the basis for $\ker A$. Thus the dual of $a_{1},\ldots,a_{n+1}$ is of the form $g_{1}=(x_{1},1),\ldots,g_{n}=(x_{n},1),g_{n+1}=(-\sum_{i=1}^{n}x_{i},-n)$. We say that the sequence $x_{1},\ldots,x_{n}$ of affinely spanning points in $\mathbb{K}^{n-d-1}$ corresponds to $a_{1},\ldots,a_{n}$. It is not difficult to see that any sequence $x_{1},\ldots,x_{n}$ of affinely spanning points in $\mathbb{K}^{n-d-1}$ corresponds to some sequence of affinely spanning points in $\mathbb{K}^{n-d-1}$. Moreover, it is an immediate consequence of Proposition 2.2 that a sequence of affinely spanning points $x_{1},\ldots,x_{n}$ in $\mathbb{K}^{n-d-1}$ which correspond to a sequence $a_{1},\ldots,a_{n}$ in $\mathbb{K}^{d}$ are in general position (i.e., no $n-d-1$ of the $x_{i}$ lie on a hyperplane in $\mathbb{K}^{n-d-1}$) if and only if $a_{1},\ldots,a_{n}$ are in general position in $\mathbb{K}^{d}$ as well. For instance, if $x_{1},\ldots,x_{n-d-1}$, say, do not affinely span $\mathbb{K}^{n-d-1}$, then $g_{1},\ldots,g_{n-d-1},g_{n}$ lie on a linear hyperplane in $\mathbb{K}^{n-d}$ and so there is an affine dependence among $a_{n-d},\ldots,a_{n}$. The proof of the converse is nearly identical.
Our condition on affinely spanning point sets $X$ with $n$ elements in $\mathbb{K}^{n-d-1}$ is essentially the Gale dual of a generic condition on affinely spanning point sets $A$ of $n$ elements in $\mathbb{K}^{d}$. We recall that a sequence $a_{1},\ldots,a_{n}$ of $n$ points in $\mathbb{K}^{d}$ is in strong general position [23] if (1) the $a_{i}$ are in ordinary general position and (2) for any non-empty pairwise disjoint subsets $I_{1},\ldots,I_{r}$ of $[n]$, one has that the (real) codimension of the intersection of the affine hulls $\operatorname{\mathrm{Aff}}(A_{j})$ of the $A_{j}=\{a_{i}\mid i\in I_{j}\}$ equals the sum of their respective codimensions, that is,
$\operatorname{\mathrm{codim}}(\cap_{j=1}^{r}\operatorname{\mathrm{Aff}}(A_{j}))=\sum_{j=1}^{r}\operatorname{\mathrm{codim}}(\operatorname{\mathrm{Aff}}(A_{j})).$
We remark that condition (1) follows from condition (2) when $\mathbb{K}=\mathbb{R}$ (see, [23]). The central property of strong general position for our purposes is that it immediately implies that any Tverberg $r$-tuple for a linear map $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$ determined by points $a_{1},\ldots,a_{n}$ in strong general position must satisfy $\sum_{j=1}^{r}|\operatorname{\mathrm{Vert}}(\sigma_{j})|\geq(r-1)(d+1)+1$.
We shall say that an affinely spanning point set $X=\{x_{1},\ldots,x_{n}\}$ in $\mathbb{K}^{n-d-1}$ is typical if $x_{1},\ldots,x_{n}$ corresponds to a sequence $a_{1},\ldots,a_{n}$ in $\mathbb{K}^{d}$ in strong general position (the $x_{i}$ are in ordinary general position and are therefore distinct). While we expect that typical sets are generic, we shall content ourselves with showing that they form an open and dense in the space of all affinely spanning point sets of size $n$ in $\mathbb{K}^{n-d-1}$. To show this, we view any sequence of $n$ points in $\mathbb{K}^{d}$ as an element of $(\mathbb{K}^{d})^{n}$, and likewise for sequences of $n$ points in $\mathbb{K}^{n-d-1}$. Let $S$ be the set of all affinely-spanning sequences of length $n$ in $(\mathbb{K}^{d})^{n}$. By generiticity of strong general position, the subset $U$ of $S$ consisting of all such sequences which are in strong general position is a dense open subset of $S$. One has the analogous situation holds for typical point sets of size $n$ in $\mathbb{K}^{n-d-1}$.
Proposition 5.1.
Let $S^{\prime}$ be the subset $(\mathbb{K}^{n-d-1})^{n}$ consisting of all affinely spanning sequences of length $n$ in $\mathbb{K}^{n-d-1}$ and let $U^{\prime}$ be the subset of $S$ consisting of all such sequences which correspond to a sequence of points in $\mathbb{K}^{d}$ in strong general position. Then $U^{\prime}$ is a dense open subset of $S^{\prime}$.
Proof of Proposition 5.1.
We first prove density. Suppose that $x=(x_{1},\ldots,x_{n})$ is a sequence of affinely spanning points in $\mathbb{K}^{n-d-1}$, so that $x$ corresponds to a sequence $a=(a_{1},\ldots,a_{n})$ of affinely spanning points in $\mathbb{K}^{d}$. Let $A$ be the corresponding matrix in the construction of Gale duality for $a_{1},\ldots,a_{n},a_{n+1}=\frac{1}{n}\sum_{i=1}^{n}a_{i}$ in $\mathbb{K}^{d}$, and let $B$ be the matrix whose columns are the conjugates of $g_{1}=(x_{1},1),\ldots,g_{n}=(x_{n},1),g_{n+1}=-\sum_{i=1}^{n}g_{i}$. By the density of $U$ in $S$, there is a sequence $(a_{m})_{m=1}^{\infty}$ of points in $(\mathbb{K}^{d})^{n}$ which are in strong general position and which converge to $a$. For each $a_{m}=(a_{1}^{m},\ldots,a_{n}^{m})$, let $a_{n+1}^{m}=\frac{1}{n}\sum_{i=1}^{n}a_{i}^{m}$, and let $A_{m}$ be the matrix whose columns are $(a_{1}^{m},1),\ldots,(a_{n+1}^{m},1)$. As the $a_{m}$ converge to $a$ and the matrices all have full rank, we have that $\ker A_{m}$ converge to $\ker A$ (i.e., if $\pi_{m}$ and $\pi$ are the projections of $\mathbb{K}^{n+1}$ onto $\ker A_{n}$ and $\ker A$, respectively, then $\pi_{m}(x)$ converges to $\pi(x)$ for all $x\in\mathbb{K}^{n+1}$). For each $m$, let $b_{1}^{m},\ldots,b_{n-d}^{m}$ be a basis for $A_{m}$ such that $b_{1}^{m},\ldots,b_{n-d-1}^{m}$ are unit vectors and $b_{n-d}^{m}=(1,\ldots,1,-n)$. Passing to a convergent subsequence if necessary when $11\leq i\leq n-d-$ 1, for each $i\in[n-d-1]$ we have that $b_{i}^{m}$ converges to some $b_{i}^{\prime}$, and that $b_{n-d}^{\prime}=(1,\ldots,1,-n)$. It is easily confirmed that $b_{1}^{\prime},\ldots,b_{n-d-1}^{\prime},b^{\prime}_{n-d}$ is a basis for $\ker A$. While the conjugates of the $b_{i}^{\prime}$ do not necessary correspond to the rows of $\overline{B}$, they do after multiplying $B$ by an appropriate invertible $(n-d)\times(n-d)$ matrix $T$. For each $m$, let $B_{m}$ be the matrix whose rows are $b_{1}^{m},\ldots,b_{n-d}^{m}$. The rows of each $TB_{m}$ are still a basis for $\ker A_{m}$, and the last row of each $TB_{m}$ is still $(1,\ldots,1,-n)$. The matrices $\overline{TB_{m}}$ converge to $\overline{B}$, and so the columns $g_{1}^{m}=(x_{1}^{m},1),\ldots,g_{n-d-1}^{m}=(x_{n-d-1}^{m},1),g_{n+1}=-\sum_{i=1}^{n}g_{i}^{m}$ of the $\overline{TB_{m}}$ converge to $g_{1},\ldots,g_{n+1}$. Thus each $x_{m}=(x_{1}^{m},\ldots,x_{n}^{m})$ lies in $U^{\prime}$ and converges to $x$.
That $U^{\prime}$ is open in $S^{\prime}$ follows similarly. For contradiction, suppose that $U^{\prime}$ were not open in $S^{\prime}$. Thus there is some $x=(x_{1},\ldots,x_{n})$ in $U^{\prime}$ corresponding to some sequence $a=(a_{1},\ldots,a_{n})$ in strong general position and a sequence $(x_{m})_{m=1}^{\infty}$ of points in $S^{\prime}\setminus U^{\prime}$ which converge to $x$. As before, let $A$ the matrix whose columns are $(a,1),\ldots,(a_{n+1},1)$, where $a_{n+1}=\frac{1}{n}\sum_{i=1}^{n}a_{i}$, and let $B$ be the matrix whose columns are $g_{1}=(x_{1},1),\ldots,g_{n}=(x_{n},1),g_{n+1}=-\sum_{i=1}^{n}g_{i}$. For each $m$, let $B_{m}$ be the matrix whose columns are $g_{1}^{m}=(x_{1}^{m},1),\ldots,g_{n}^{m}=(x_{n}^{m},1),g^{m}_{n+1}=-\sum_{i=1}^{n}g^{m}_{i}$, and choose a basis $b_{1}^{m},\ldots,b_{n-d}^{m}$ for $\ker B_{m}$ so that $b_{n-d}$ is the all 1 vector and the other $b_{i}^{m}$ have norm 1. As $\ker B_{m}$ converges to $\ker B$, after passing to a subsequence if necessary we may assume that each $b_{i}^{m}$ converges to some $b_{i}^{\prime}$, where $b_{n-d}^{\prime}$ is the all one vector. Again, it is easily confirmed that the vectors $b_{1}^{\prime},\ldots,b_{n-d}^{\prime}$ are a basis for $\ker B$. After applying an appropriate invertible transformation to the $b_{i}^{\prime}$ and applying this transform to $b_{1}^{m},\ldots,b_{n-d-1}^{m}$ for all $m$, we may assume that $b_{i}^{\prime}=b_{i}$ for all $1\leq i\leq n-d$. We now let $A_{m}$ be the matrix whose rows are the conjugates of the $b_{i}^{m}$, and we let $a_{m}=(a_{1}^{m},\ldots,a_{n}^{m})$ be its first $n$ columns excluding the final row of ones. We have $a_{n+1}^{m}=\frac{1}{n}\sum_{i=1}^{n}a_{i}$ for all $m$ by construction. As the rows of $A$ are the conjugates of the $b_{i}$, $A_{m}$ converges to $A$ and therefore $a_{m}=(a_{1}^{m},\ldots,a_{n}^{m})$ converges to $a$. On the other hand, each $x_{m}$ corresponds to $a_{m}$, and since $x_{m}$ is in $S^{\prime}\setminus U^{\prime}$, each $a_{m}$ is in $S\setminus U$, thereby contradicting that $U$ is open in $S$. ∎
We now show that the open half-flats of the $r$-fans of Theorems 1.4 and  1.6 must share a large fraction of the points of any typical point set.
Proposition 5.2.
Let $d,m\geq 1$ be integers and let $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$. Let $n\geq(r-1)(d+m+1)+1$ if $\mathbb{K}=\mathbb{R}$, let $n\geq(r-1)(2d+m+1)+1$ if $\mathbb{K}=\mathbb{C}$, and suppose that $X$ is a typical set of $n$ points in $\mathbb{K}^{n-d-1}$. Then the $r$-fans $F_{r}=\cup_{j=1}^{r}B_{j}$ guaranteed by Theorem  1.4 and 1.6 satisfy the property that $\sqcup_{j=1}^{r}\operatorname{\mathrm{Int}}(B_{j})$ contains at least $(r-1)(d+1)+1$ points of $X$ when $\mathbb{K}=\mathbb{R}$, and at least $(r-1)(2d+1)+1$ points of $X$ when $\mathbb{K}=\mathbb{C}$.
Proof of Proposition 5.2.
Suppose that $X=\{x_{1},\ldots,x_{n}\}$ is a typical point set in $\mathbb{K}^{n-d-1}$. Thus $x_{1},\ldots,x_{n}$ corresponds to a sequence of points $a_{1},\ldots,a_{n}$ in $\mathbb{K}^{d}$ in strong general position. The distributing $r$-fan $F_{r}=\cup_{j=1}^{r}$ guaranteed by the proofs of Theorem 1.4 and Theorem 1.6 was given by a Tverberg $r$-tuple $(\sigma_{1},\ldots,\sigma_{r})$ for the linear map $f\colon\Delta_{n-1}\rightarrow\mathbb{K}^{d}$ determined by the $a_{i}$. Moreover, we had $|\operatorname{\mathrm{Int}}(B_{j})\cap X|=|\operatorname{\mathrm{Vert}}(\sigma_{j})|$ for all $j\in[r]$. As the $a_{i}$ are in strong general position, we have $\sum_{j=1}^{r}|\operatorname{\mathrm{Int}}(B_{j})\cap X|=\sum_{j=1}^{r}|\operatorname{\mathrm{Vert}}(\sigma_{j})|\geq(r-1)(d+1)+1$. ∎
5.2. Asymptotics of Theorem 1.4
We conclude this section by proving the near optimality of the real case of Theorem 1.4 with respect to dimension for all prime powers $r\geq 3$ when $m\geq 2$, and likewise in the complex setting when $r=2$ and $m\geq 2$. For this, it suffices to consider the near optimality of Theorems 1.1 and  1.3, respectively.
Proposition 5.3.
Let $d\geq 1$ and $m\geq 2$ be integers.
  1. (a)
    Let $r\geq 3$, let $k\in\{0,\ldots,r-1\}$, and let $n=(r-1)(d+m+1)+k+1$. If $\ell>2+\frac{k}{r-1}$, then there exists a set $X$ of $n+\ell$ affinely spanning points in $\mathbb{R}^{n-d-1}$ and an $m$-coloring of $X$ which cannot be equidistributed by any $r$-fan of codimension $r-2$.
  2. (b)
    Let $r=2$, let $k\in\{0,1\}$, and let $n=2(d+1)+m+k$. If $\ell>2+\frac{k}{2}$, then there exists a set $X$ of $n+\ell$ complex affinely spanning points in $\mathbb{C}^{n-d-1}$ and an $m$-coloring of $X$ which cannot be equdistributed by any complex regular $2$-fan in $\mathbb{C}^{n-d-1}$.
Proof of Proposition 5.3.
In the real case, let $n=(r-1)(d+m+1)+k+1$, where $10\leq k\leq r-$ 1, and assume that $\ell>2+\frac{k}{r-1}$. Suppose that $a_{1},\ldots,a_{n},a_{n+\ell-1}$ are points in strong general position in $\mathbb{R}^{d+\ell-1}$, and let $x_{1},\ldots,x_{n+\ell-1}$ be their Gale dual in $\mathbb{R}^{n-d-1}$. These points are all in general linear position in $\mathbb{R}^{n-d-1}$ (i.e., any $n-d-1$ of them linearly span $\mathbb{R}^{n-d-1}$), so in particular none of the $x_{i}$ is zero. Now consider the sequence $a_{1}^{\prime}=(a_{,}1),\ldots,a_{n+\ell-1}^{\prime}=(a_{n+\ell-1},1)$ in $\mathbb{R}^{d+\ell}$, and append to this sequence any vector $a_{n+\ell}^{\prime}$ in $\mathbb{R}^{d+\ell}$ such that $a_{1}^{\prime},\ldots,a^{\prime}_{n+\ell}$ affinely span $\mathbb{R}^{d+\ell}$. We now claim that, for a suitable choice of basis, the Gale transform of the $a_{i}^{\prime}$ is $x_{1},\ldots,x_{n+\ell-1},x_{n+\ell}$, where $x_{n+\ell}=0$. To see this, let $A$ be the $(d+\ell)\times(n+\ell-1)$ matrix in the construction of the Gale transform of $a_{1},\ldots,a_{n+\ell-1}$, and let $b_{1},\ldots,b_{n-d-1}\in\mathbb{R}^{n+\ell-1}$ be a choice of basis for $\ker A$ so that $x_{1},\ldots,x_{n+\ell-1}$ are the columns of the matrix $B$ whose rows are the $b_{i}$. Letting $A^{\prime}$ be the matrix in the construction of the Gale transform of the $a_{i}^{\prime}\in\mathbb{R}^{d+\ell}$, we may choose $b_{1}^{\prime}=(b_{1},0),\ldots,b_{n-d-1}^{\prime}=(b_{n-d-1},0)$ as a basis for $\ker A^{\prime}$. Thus the Gale transform of $a_{1}^{\prime},\ldots,a^{\prime}_{n+\ell}$ in $\mathbb{R}^{n+\ell}$ is $x_{1},\ldots,x_{n+\ell-1},x_{n+\ell}=0$ as claimed. As any $n-d-1$ of $x_{1},\ldots,x_{n+\ell-1}$ linearly span $\mathbb{R}^{n-d-1}$ and $x_{1},\ldots,x_{n+\ell-1}$ sums to zero, it follows easily that $x_{1},\ldots,x_{n+\ell}$ affinely span $\mathbb{R}^{n-d-1}$.
We now show that there exists an $m$-coloring of $X=\{x_{1},\ldots,x_{n+\ell}\}$ which cannot be equidstributed by any $r$-fan in $\mathbb{R}^{n-d-1}$ of codimension $r-2$. Namely, let $X=X_{1}\sqcup\cdots\sqcup X_{m}$ be an $m$-coloring such that $|X_{i}|=r-1$ for all $i\geq 2$ and $x_{n+\ell}=0$ is in $X_{m}$. Now suppose that $F_{r}=\cup_{j=1}^{r}B_{j}$ is an $r$-fan of codimension $r-2$ which equidistributes this coloring. As $|\operatorname{\mathrm{Int}}(B_{j})\cap X_{i}|\leq|X_{i}|/r$ for all $i\in[m]$ and $j\in[r]$, all the points of $X\setminus X_{1}$ must lie on the center $C$ of $F_{r}$. In particular, $F_{r}$ is linear. Now let $f\colon\Delta_{n+\ell-1}\rightarrow\mathbb{R}^{d+\ell}$ be the linear map determined by the $a_{i}^{\prime}$ and let $[n+\ell]=C_{1}\sqcup\cdots\sqcup C_{m}$ be the $m$-coloring of the vertex set of $\Delta_{n+\ell-1}$ corresponding to the $m$-coloring of $X$. Applying part (b) of Lemma 3.1, there is a proper Tverberg $r$-tuple $(\sigma_{1},\ldots,\sigma_{r})$ for $f$ such that each $I_{j}:=\operatorname{\mathrm{Vert}}(\sigma_{j})$ is in bijective correspondence with $X\cap\operatorname{\mathrm{Int}}(B_{j})$ and $[n+\ell-1]\setminus\sqcup_{j=1}^{r}I_{j}$ is in bijective correspondence with $C\cap X$. Thus $|C_{i}\cap\operatorname{\mathrm{Vert}}(\sigma)|=|X_{i}\cap\operatorname{\mathrm{Int}}(B_{j})|$ for all $j\in[r]$ and all $i\in[m]$. Thus each $I_{j}$ is contained in the color class $C_{1}\subset[n+\ell-1]$, which consists of $n-(m-1)(r-1)=(r-1)(d+2)+k+1$ points. Letting $\Delta_{C_{1}}$ be the simplex generated by $C_{1}$, we have that the image of the restriction $f\colon\Delta_{C_{1}}\rightarrow\mathbb{R}^{d+\ell}$ of $f$ has image in $\mathbb{R}^{d+\ell-1}\times\{1\}$. Since the points $a_{1}^{\prime}=(a_{1},1),\ldots,a_{n+\ell-1}^{\prime}=(a_{n+\ell-1},1)$ are in strong general position in $\mathbb{R}^{d+\ell-1}\times\{1\}$ and $k<(r-1)(\ell-2)$, we have
$\dim(\cap_{j=1}^{r}\operatorname{\mathrm{Aff}}(f(\sigma_{j})))=\sum_{j=1}^{r}\operatorname{\mathrm{codim}}(\operatorname{\mathrm{Aff}}(f(\sigma_{j}))\geq r(d+\ell)-[(r-1)(d+2)+k+1]>d+\ell-1,$
so $\cap_{j=1}^{r}\operatorname{\mathrm{Aff}}(f(\sigma_{j}))=\emptyset$ and therefore $(\sigma_{1},\ldots,\sigma_{r})$ is not a Tverberg $r$-tuple for $f$. Thus there is no $r$-fan of codimension $r-2$ which equidistributes the $m$-coloring of $X$.
The proof in the complex case when $r=2$ is analogous. Let $n=2(d+1)+m+k$ where $k\in\{0,1\}$. As before, there exist $n+\ell$ complex affinely spanning points $a_{1},\ldots,a_{n+\ell}$ in $\mathbb{C}^{d+\ell}$ whose Gale dual $x_{1},\ldots,x_{n+\ell}$ in $\mathbb{C}^{n-d-1}$ satisfies $x_{n+\ell}=0$ and for which $a_{1},\ldots,a_{n+\ell-1}$ are in strong general position in $\mathbb{C}^{d+\ell-1}\times\{1\}$. In particular, $X$ complex affinely spans $\mathbb{C}^{n-d-1}$. As before, we choose an $m$-coloring of $X=\{x_{1},\ldots,x_{n+\ell}\}$ so that $|X_{i}|=1$ for all $2\leq i\leq m$ and $0\in X_{m}$. If $F_{2}$ were a complex regular $2$-fan which equidistributes the $m$-coloring, then $F_{2}$ would be linear and so by Remark 3.3 the linear map $f\colon\Delta_{n+\ell-1}\rightarrow\mathbb{C}^{d+\ell}$ determined by the $a_{i}$ would have a Tverberg pair $(\sigma_{1},\sigma_{2})$ where each $\sigma_{i}$ does not contain $n+\ell$. As $a_{1},\ldots,a_{n+\ell-1}$ are in strong general position in $\mathbb{C}^{d+\ell-1}\times\{1\}$, as before the condition that $k<2(\ell-2)$ gives $\operatorname{\mathrm{codim}}(\operatorname{\mathrm{Aff}}(f(\sigma_{1})\cap\operatorname{\mathrm{Aff}}(f(\sigma_{2}))>2(d+\ell-1)$. We conclude as before that no equidistributing complex regular $2$-fan exists. ∎
Corollary 1.2 is now an immediate consequence of Theorem 1.1 and Proposition 5.3.
Proof of Corollary 1.2.
Let $c=(r-1)(m+1)$ and let $d=(r-2)s+t+c$ where $s\geq 1$ and $t\in\{0,\ldots,r-2\}$. Letting $n=d+s+1$ and applying Theorem 1.1, we have that any $m$-coloring of any affinely spanning set of $n=(r-1)(s+m+1)+t+1$ points in $\mathbb{R}^{n-s-1}=\mathbb{R}^{d}$ can be equidistributed by an $r$-fan in $\mathbb{R}^{d}$ of codimension $r-2$. Thus $n(r,m,d)\geq n$. On the other hand, Proposition 5.3 shows that $n(r,m,d)\leq n+\ell-1$ so long as $\ell>2+\frac{t}{r-1}$. But $20\leq t\leq r-$ 2, so $n(r,m,d)\leq n+2$. ∎
Remark 5.4.
As in the real case, we may denote by $n_{\mathbb{C}}(r,m,d)$ the maximum number $n$ such that any $m$-coloring of any set of $n$ complex affinely spanning points in $\mathbb{C}^{d}$ can be equidistributed by a complex regular $r$-fan in $\mathbb{C}^{d}$. Letting $d=s+t+m+1$ with $s\geq 1$ and $t\in\{0,1\}$ and setting $n=d+s+1$, Theorem 1.3 and Proposition 5.3 likewise give the bound $n\leq n_{\mathbb{C}}(2,m,d)\leq n+2$ for complex regular $2$-fans.
6. A Sarkaria-type Theorem for Two Tverberg Pairs
Our equidistribution results for two $r$-fans follow from an extension of Sarkaria’s theorem to two Tverberg $r$-tuples. Our proof will follow the Configuration Space–Test Map paradigm typically used in topological combinatorics, and in particular will closely match that of an extension of the $r=2$ case of Sarkaria’s theorem given in [12]. In our case, the desired pair of Tverberg tuples for the given map $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$ will be shown to be equivalent to the existence of a zero of a naturally associated equivariant map from a certain topological space equipped with a $(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$-action to an associated $(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$-module. The proof is completed by a Borsuk-Ulam type statement (Proposition 6.1) which ensures this map has a zero.
6.1. Configuration Space
Our configuration space is based on the “deleted product” scheme commonly used in topological Tverberg-type theory (see, e.g., [22, 18, 33]). Given a simplex $\Delta_{n-1}$, recall that the support $\mathrm{supp}(x)$ of a point $x$ in $\Delta_{n-1}$ is the smallest (closed) face in $\Delta_{n-1}$ containing $x$. For $r\geq 2$, the deleted $r$-fold product
$(\Delta_{n-1})^{\times r}_{\Delta}=\{x=(x_{0},\ldots,x_{r-1})\mid x_{i}\in\Delta_{n-1}\,\,\text{for all}\,\,i\,\,\text{and}\,\,\mathrm{supp}(x_{i})\cap\mathrm{supp}(x_{j})=\emptyset\,\,\text{for all}\,\,i\neq j\}$
is the subset of the Cartesian product $\Delta_{n-1}^{\times r}$ consisting of all $r$-tuples of elements of $\Delta_{n-1}$ whose supports are pairwise disjoint. There is a natural free $\mathbb{Z}_{r}$-action on $(\Delta_{n-1})^{\times r}_{\Delta}$, so that for each $i\in\mathbb{Z}_{r}$ one cyclically shifts each coordinate of an $r$-tuple $x=(x_{0},\ldots,x_{r-1}$ forward by $i$. We shall be concerned with the Cartesian product
$X(r,n)=(\Delta_{n-1})^{\times r}_{\Delta}\times(\Delta_{n-1})^{\times r}_{\Delta}$
of the deleted product with itself, which we equip the standard $(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$-action on $X(r,n)$. Thus each $\mathbb{Z}_{r}$-factor of $\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}$ acts independently on the corresponding $(\Delta_{n-1})^{\times r}_{\Delta}$ factor of $X(r,n)$.
6.2. Test Space
Our test space will be the direct sum of two $(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$-modules. First, consider the regular representation $\mathbb{R}[\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}]=\{\sum_{(i,j)\in\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}}r_{(i,j)}\cdot(i,j)\mid r_{(i,j)}\in\mathbb{R}\,\,\text{for all}\,\,(i,j)\in\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}\}$, for which the $(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$-action is given by setting $(k,\ell)\cdot(i,j)=(k+i,\ell+j)$ for each $(k,\ell),(i,j)\in\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}$ and extending linearly. Inside the regular reperesentation is the subrepresentation $U_{r}=\{\sum_{(i,j)\in\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}}r_{(i,j)}\cdot(i,j)\mid\sum_{(i,j)}r_{i,j}=0\}$ consisting of all elements of the regular representation whose coefficients sum to zero. In other words, $U_{r}$ is the complement of the diagonal 1-dimensional trivial subrepresentation in $\mathbb{R}[\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}]$ and so has dimension $r^{2}-1$. For each $m\geq 1$, we then let $U_{r}^{\oplus m}$ be the $m$-fold direct sum of $U_{r}$, which we equip with the diagonal $(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$-action.
For the second representation, we consider $\mathbb{R}^{d}[\mathbb{Z}_{r}]=\{\sum_{i\in\mathbb{Z}_{r}}y_{i}\cdot i\mid y_{i}\in\mathbb{R}^{d}\}$, which can be identified with the $d$-fold sum of the regular representation of $\mathbb{Z}_{r}$. As before, this is defined by setting $i\cdot j=i+j$ for each $i,j\in\mathbb{Z}_{r}$ and extending linearly, which results in a forward cyclic shift of the coefficients. We now let let $V_{d}=\{\sum_{i\in\mathbb{Z}_{r}}y_{i}\cdot i\mid\sum_{i}y_{i}=0\}$ be the orthogonal complement of the diagonal $d$-dimensional trivial subrepresentation inside $\mathbb{R}^{d}[\mathbb{Z}_{r}]$. Letting each $\mathbb{Z}_{r}$-factor of $\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}$ act independently on each $V_{d}$ factor of $V_{d}\oplus V_{d}$ defines a $(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$-subrepresentation of $\mathbb{R}[\mathbb{Z}_{r}]\oplus\mathbb{R}[\mathbb{Z}_{r}]$ of dimension $2d(r-1)$.
Finally, we let
$U:=(V_{d}\oplus V_{d})\oplus U_{r}^{\oplus m}.$ (6.1)
6.3. Test Map
Let $d\geq 1$ be and integer, let $r\geq 3$ be an odd integer, and suppose that $n\geq(r-1)(d+1)+m\frac{r^{2}-1}{2}+1$. We assume that $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$ is a continuous map and that $\mathcal{F}$ is a family of non-empty subsets of $[n]=\operatorname{\mathrm{Vert}}(\Delta_{n-1})$ such that $\chi(\mathrm{KG}^{r^{2}}\mathcal{F})\leq m$. Thus $\mathcal{F}=\mathcal{F}_{1}\cup\cdots\cup\mathcal{F}_{m}$, where none of the $m$ subfamilies $F_{k}$ contains $r^{2}$ pairwise disjoint sets. Without loss of generality, we may assume that each $\mathcal{F}_{k}$ is closed under taking supersets.
First, we construct a continuous map $T\colon X(r,n)\rightarrow V_{d}\oplus V_{d}$ whose zeros correspond to a pair of Tverberg tuples. For each $x=(x_{0},\ldots,x_{r-1})\in(\Delta_{n-1})^{\times r}_{\Delta}$, we let $\mathrm{Avg}(f,x)=\frac{1}{r}\sum_{i=0}^{r-1}f(x_{i})$ denote the average value of the $f(x_{i})$, and for each $(x,y)\in X(r,n)$ we set
$T(x,y)=(\sum_{i\in\mathbb{Z}_{r}}(f(x_{i})-\mathrm{Avg}(f,x))\cdot i)\oplus(\sum_{i\in\mathbb{Z}_{r}}(f(y_{i})-\mathrm{Avg}(f,y)\cdot i)).$
Clearly, both $(\mathrm{supp}(x_{0}),\ldots,\mathrm{supp}(x_{r-1}))$ and $(\mathrm{supp}(y_{0}),\ldots,\mathrm{supp}(y_{r-1}))$ are Tverberg tuples for $f$ if and only if $T(x,y)=0$. It is easily confirmed that $T$ is $(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$-equivariant with respect to the actions on the domain and codomain.
Next, we construct a continuous map $D\colon X(r,n)\rightarrow\mathbb{R}^{\perp}[\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}]^{\oplus m}$ which ensures that, for each zero $(x,y)$ of $T$, none of the $r^{2}$ subsets of $[n]$ of the form $\operatorname{\mathrm{Vert}}(\mathrm{supp}(x_{i})\cap\mathrm{supp}(y_{j}))$ contains any element from $\mathcal{F}$. As each $\mathcal{F}_{k}$ is closed under taking supersets, $\Sigma_{k}:=\{\sigma\mid\sigma\notin\mathcal{F}_{k}\}$ is an abstract simplicial complex on $[n]$, namely the one defined by the rule that its set of minimal non-faces is $\mathcal{F}_{k}$. In particular, any $\operatorname{\mathrm{Vert}}(\mathrm{supp}(x_{i})\cap\mathrm{supp}(y_{j}))$ is an element of $\Sigma_{k}$ if and only if it does not contain any elements of $\mathcal{F}_{k}$. We denote by $\|\Sigma_{k}\|$ the realization of $\Sigma_{k}$ as a geometric subcomplex of $\Delta_{n-1}$.
Endowing $\Delta_{n-1}\times\Delta_{n-1}$ with the product metric $d\colon\Delta_{n-1}\times\Delta_{n-1}\rightarrow[0,\infty)$ (say, with the Euclidian metric on each $\Delta_{n-1}$ factor), for each $k\in[m]$ we define the map $d_{k}\colon X(r,n)\rightarrow U_{m}$ by setting
$d_{k}(x,y)=\sum_{(i,j)\in(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})}\left[\left(d((x_{i},y_{j}),\|\Sigma_{k}\|\times|\Sigma_{k}\|\right)-a(x,y)\right]\cdot(i,j)\in\mathbb{R}[\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}],$
where $a(x,y)=\sum_{(i,j)\in(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})}\left(d(x_{i},y_{j}),\|\Sigma_{k}\|\times|\Sigma_{k}\|\right)$ is the average of the distances $d((x_{i},y_{j}),\|\Sigma_{k}\|\times\|\Sigma_{k}\|)$ from $(x_{i},y_{j})$ to $\|\Sigma_{k}\|\times\|\Sigma_{k}\|$. Again, it is easily verified that each $d_{k}$ is equivariant with respect to the actions concerned. We note that $d((x_{i},y_{j}),\|\Sigma_{k}\|\times\|\Sigma_{k}\|)=0$ if and only if both $x_{i}$ and $y_{j}$ lie in $\|\Sigma_{k}\|$, which is the case if and only $\mathrm{supp}(x_{i})\cap\mathrm{supp}(y_{j})$ lies in $\|\Sigma_{k}\|$, or in other words that $\operatorname{\mathrm{Vert}}(\mathrm{supp}(x_{i})\cap(\mathrm{supp}(y_{j}))$ does not contain any element of $\mathcal{F}_{k}$ as a subset. We then define
$D=(d_{1},\ldots,d_{m})\colon X(r,n)\rightarrow U_{k}^{\oplus m}.$
For each $k\in[m]$, we have that $d_{k}(x,y)=0$ if and only if the distance from each $(x_{i},y_{j})$ to $\|\Sigma_{k}\|\times\|\Sigma_{k}\|$ is equalized. On the other hand, for each $(x,y)\in X(r,n)$ the $r^{2}$ intersections $\mathrm{supp}(x_{i})\cap\mathrm{supp}(y_{i})$ are pairwise disjoint, and since $\mathcal{F}_{k}$ does not contain $r^{2}$ pairwise disjoint sets there must be some $(i_{0},j_{0})\in\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}$ such that $\operatorname{\mathrm{Vert}}(\mathrm{supp}(x_{i_{0}})\cap(\mathrm{supp}(y_{j_{0}}))\in\Sigma_{k}$. Thus $d(x_{i_{0}},y_{j_{0}}),\|\Sigma_{k}\|\times\|\Sigma_{k}\|)=0$. We conclude therefore that $d_{k}(x,y)=0$ if and only if each $\operatorname{\mathrm{Vert}}(\mathrm{supp}(x_{i})\cap(\mathrm{supp}(y_{j}))$ does not contain any element of $\Sigma_{k}$.
Finally, we let
$F\colon=T\oplus D\colon X(r,n)\rightarrow U,$ (6.2)
which is again $(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$-equivariant. Given the discussion above, we have that $F(x,y)=0$ if and only if $(\mathrm{supp}(x_{0}),\ldots,\mathrm{supp}(x_{r-1}))$ and $(\mathrm{supp}(y_{0}),\ldots,\mathrm{supp}(y_{r-1}))$ are Tverberg pairs for $f$ such that no element of $\mathcal{F}$ is a subset of any $\operatorname{\mathrm{Vert}}(\mathrm{supp}(x_{i})\cap\mathrm{supp}(y_{j}))$. Theorem 1.8 then is then an immediate consequence of the following proposition, whose proof we defer to the next section.
Proposition 6.1.
Let $d\geq 1$ be an integer, let $r\geq 3$ be an odd prime, let $n\geq(r-1)(d+1)+m\frac{r^{2}-1}{2}$ be an integer, and let $U$ be the representation (6.1). If the $r$-ary expansion of $m(r-1)/2$ has all even coefficients, then any continuous $(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$-equivariant map $F\colon X(r,n)\rightarrow U$ has a zero.
Proof of Theorem 1.8.
Let $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$ be continuous, and suppose that $\mathcal{F}$ is a non-empty family of subsets of $[n]$ for which $\chi(\mathrm{KG}^{r^{2}}(\mathcal{F}))\leq m$. As the map $F\colon X(r,n)\rightarrow U$ given by (6.2) is equivariant, by Proposition 6.1 there is some $(x,y)\in X(r,n)$ such that $F(x,y)=0$. Hence $(\mathrm{supp}(x_{0}),\ldots,\mathrm{supp}(x_{r-1}))$ and $(\mathrm{supp}(y_{0}),\ldots,\mathrm{supp}(y_{r-1}))$ are the desired Tverberg pairs for $f$.∎
Remark 6.2.
We conclude this section by verifying that the particular values of $m$ mentioned in the introduction apply to Theorem 1.8 (as well as to Theorem 1.7 and Theorem 8.1 below).
First, suppose that $m=2a(r^{\ell_{1}}+\cdots+r^{\ell_{k}})$ where $\ell_{1}<\cdots<\ell_{k}$ and $11\leq a\leq r-$ 1 is odd. We have
$m(r-1)/2$ $=(r-1)a(r^{\ell_{1}}+\cdots+r^{\ell_{k}})$
$=(r-a)r^{\ell_{1}}+(a-1)r^{\ell_{1}+1}+(r-a)r^{\ell_{2}}+(a-1)r^{\ell_{2}+1}\cdots+(r-a)r^{\ell_{k}}+(a-1)r^{\ell_{k}+1}.$
As both $a-1$ and $r-a$ are even, we are done provided that $\ell_{i}+1<\ell_{i+1}$ for all $i\in[k]$. On the other hand, if $\ell_{i}=\ell_{i+1}$ for any $i$, then the coefficient of $r^{\ell_{i+1}}$ in the $r$-ary expansion of $m(r-1)/2$ is $r-1$, which is again even. Thus each coefficient in the $r$-ary expansion of $m(r-1)/2$ is even. As $1+r+\cdots+r^{\ell}=\frac{r^{\ell+1}-1}{r-1}$, Theorem 1.8 applies in particular if $m=2a(r^{\ell+1}-1)/(r-1)$ where $\ell\geq 0$.
7. Proof of Proposition 6.1
Our proof of Proposition 6.1 follows a standard factorization trick, whereby the given equivariant map induces a section of a complex vector bundle given by the Borel construction. One then shows that the bundle’s top Chern class is non-zero, so that the section, and therefore the equivariant map, necessarily has a zero. A similar Chern class computation can be found in  [30] in the context of mass equipartitions by complex regular fans. We refer the reader to  [21, 17] for the basic theory of vector bundles and characteristic classes.
7.1. Complex Representations
We begin by showing that the real $(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$-module $U$ (6.1) above can be equivariantly identified with the complex $(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$-module $W$ (7.1) below.
To that end, recall that as $(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$ is a finite abelian group, its irreducible complex representations are all 1-dimensional and correspond to the elements of the group itself. Explicitly, letting $\omega_{r}=\exp(2\pi i/r)$ be the standard $r$-th root of unity, each $(i,j)\in\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}$ gives rise to the 1-dimensional module $W_{(i,j)}=\mathbb{C}$ on which the group acts by setting $(k,\ell)\cdot z=\omega_{r}^{ki+\ell j}z$ for each $z\in\mathbb{C}$ and each $(k,\ell)\in(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$. Next, recall that the complexification $V\otimes\mathbb{C}$ of any finite dimensional real $(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$-representation $V$ results in a complex $(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$-representation of the same dimension. For instance, the complexification of the real regular representation $\mathbb{R}[(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})]$ is the complex regular representation $\mathbb{C}[\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}]=\{\sum_{(i,j)\in\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}}z_{(i,j)}\cdot(i,j)\mid z_{(i,j)}\in\mathbb{C}\}$, and likewise the complexifications of $V_{d}\oplus V_{d}$ or $U_{r}$ of Section 6 are obtained by replacing real with complex coefficients in their respective definitions. One then has the decompositions $(V_{d}\oplus V_{d})\otimes\mathbb{C}\cong\oplus_{1\leq i,j\leq r-1}(W_{i,0}^{\oplus d}\oplus W_{0,j}^{\oplus d})$ and $U_{r}^{\oplus m}\otimes\mathbb{C}\cong\oplus_{(i,j)\neq 0}W_{i,j}^{\oplus m}$ into complex irreducible representations.
On the other hand, for any complex $(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$-representation $V$ one may consider the $2d$-dimensional real $(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$-representation $V_{\mathbb{R}}$ given by restricting scalar multiplication in $V$ to $\mathbb{R}$. It follows easily (see, e.g.,  [13]) that $(U\otimes\mathbb{C})_{\mathbb{R}}\cong U\oplus U$ for any real representation $U$. Moreover, by identifying $\omega_{r}^{k}$ with the corresponding real $2\times 2$ rotation matrix, one sees that $(W_{i,j})_{\mathbb{R}}$ and $(W_{-i,-j})_{\mathbb{R}}$ are isomorphic for any $(i,j)\in(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$. Now let $A$ be the subset of $\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}$ consisting of all non-zero $(i,j)$ whose last non-zero coordinate lies in $\{1,\ldots,\frac{r-1}{2}\}$. One then has that
$(U\oplus U)\cong\oplus_{1\leq i,j\leq\frac{r-1}{2}}(W_{(i,0))})_{\mathbb{R}}^{\oplus 2d}\oplus(W_{(0,j)})_{\mathbb{R}}^{\oplus 2d})\oplus\oplus_{(i,j)\in A}(W_{(i,j)})_{\mathbb{R}}^{\oplus 2m}.$
Thus the real representation $U$ may be equivariantly identified with complex representation
$W:=\oplus_{1\leq i,j\leq\frac{r-1}{2}}(W_{(i,0)}^{\oplus d}\oplus W_{(0,j)}^{\oplus d})\oplus\oplus_{(i,j)\in A}W_{(i,j)}^{\oplus m}.$ (7.1)
In particular, $\dim W=N:=(r-1)d+m\frac{r^{2}-1}{2}.$
7.2. A Chern Class Computation
Let $n>r$ be an integer. For simplicity, we will let $X_{n}=(\Delta_{n-1})^{\times r}_{\Delta}$ denote the deleted $r$-fold product of Section 6 and denote by $\overline{X_{n}}=X_{n}/\mathbb{Z}_{r}$ its quotient by the free $\mathbb{Z}_{r}$-action discussed there. The quotient of $X(r,n)=X_{n}\times X_{n}$ under the standard $(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$-action is then $\overline{X(r,n)}=\overline{X_{n}}\times\overline{X_{n}}$.
For our purposes, the central feature of $X_{n}$ is that it is an $(n-r)$-dimensional cellular complex which is $(n-r-1)$-connected (see, e.g, [4]). The union $\bigcup_{n=r}^{\infty}X_{n}$ may therefore be taken to be the total space $E\mathbb{Z}_{r}$ of the classifying bundle $\mathbb{Z}_{r}\hookrightarrow E\mathbb{Z}_{r}\rightarrow B\mathbb{Z}_{r}:=E\mathbb{Z}_{r}/\mathbb{Z}_{r}$ of the group $\mathbb{Z}_{r}$. The classifying space $B\mathbb{Z}_{r}$ may then be taken to be $\bigcup_{n=r}^{\infty}\overline{X_{n}}$, so that each $\overline{X_{n}}$ is the $(n-r)$-dimensional skeleton of $B\mathbb{Z}_{r}$. We may then take $B(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})=B\mathbb{Z}_{r}\times B\mathbb{Z}_{r}$ to be the classifying space of $\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}$, of which $\overline{X(r,n)}$ is then a subcomplex.
Now let $n\geq(r-1)(d+1)+m\frac{r^{2}-1}{2}+1$ and suppose that $F\colon X(r,n)\rightarrow W$ is a $G$-equivariant map. As above, let $N=(r-1)d+m(r^{2}-1)/2$. Corresponding to the $N$-dimensional representation $W$, let $E(r,n):=(X(r,n))\times_{\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}}W$ be the $N$-dimensional complex vector bundle over $\overline{X(r,n)}=X(r,n)$ obtained by quotienting the trivial bundle $X(r,n)\times W$ under the diagonal $(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$-action. The map $F$ gives rise to a section $s\colon\overline{X(r,n)}\rightarrow E(r,n)$ of this bundle which is induced by the equivariant section $X(r,n)\rightarrow X(r,n)\times W$ of the trivial bundle given by $x\mapsto(x,F(x))$. Thus $F$ has a zero if and only if the section $s$ does. On the other hand, the existence of a never-vanishing section of $E(r,n)$ implies that the bundle’s top Chern class $c_{N}(E(r,n))\in H^{2N}(\overline{X(r,n)};\mathbb{Z})$ is zero. Thus the proof will be complete once it is shown that $c_{N}(E(r,n))\neq 0$.
Recall that the total Chern class $c(V)$ of any finite-dimensional complex representation $V$ of a finite group $G$ is defined to be the total Chern class of the vector bundle $E_{V}:=EG\times_{G}V$ over its classifying space $BG$ (see, e.g., [1]). In our case, it is easily seen that the bundle $E(r,n)$ above is the pullback under inclusion $i\colon X(r,n)\hookrightarrow B(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$ of the bundle $E_{W}$ determined by the representation $W$ (7.1). Thus $E(r,n)=i^{*}(c_{N}(W))$, and so $c_{N}(E(r,n))=i^{\ast}(c_{N}(W))$ by the naturality of Chern classes.
We now calculate $c_{N}(W)\in H^{\ast}(B(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r});\mathbb{Z})$. To that end, the Künneth formula (see, e.g., [16]) shows that the map
$H^{\ast}(B\mathbb{Z}_{r};\mathbb{Z})\otimes H^{\ast}(B\mathbb{Z}_{r};\mathbb{Z})\stackrel{{\scriptstyle\pi_{1}^{\ast}\otimes\pi_{2}^{\ast}}}{{\longrightarrow}}H^{*}(B(\mathbb{Z}_{r}\oplus);\mathbb{Z})$
induced by the projection maps of $B(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})=B\mathbb{Z}_{r}\times B\mathbb{Z}_{r}$ onto its $B\mathbb{Z}_{r}$ factors is an injection. We now show that $c_{N}(W)$ lies in the image of $H^{\ast}(B\mathbb{Z}_{r};\mathbb{Z})\otimes H^{\ast}(B\mathbb{Z}_{r};\mathbb{Z})$ under this map, which we call the tensor subring $H^{\ast}_{\operatorname{\mathrm{tensor}}}(B(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r});\mathbb{Z})$ of $H^{\ast}(B(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r});\mathbb{Z})$. To see this, first recall that $H^{\ast}(B\mathbb{Z}_{r})=\mathbb{Z}[y]/(ry)$, where $y=c_{1}(W_{1})\in H^{2}(B\mathbb{Z}_{r};\mathbb{Z})\cong\mathbb{Z}_{r}$ can be taken to be the total Chern class of the standard 1-dimensional representation $W_{1}$ of $\mathbb{Z}_{r}$ (see, e.g., [1] or  [16]) arising from the linear $\mathbb{Z}_{r}$-action on $\mathbb{C}$ given by $i\cdot z=\omega_{r}^{i}z$ for all $i\in\mathbb{Z}_{r}$ and $z\in\mathbb{C}$. On the other hand, $E_{W_{(1,0)}}=\pi_{1}^{\ast}(E_{W_{1}})$ and $E_{W_{(0,1)}}=\pi_{2}^{\ast}(E_{W_{1}})$ are the pullbacks of $E_{W_{1}}$ under the respective coordinate projections $\pi_{1}$ and $\pi_{2}$ of $B(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$ onto its respective $B\mathbb{Z}_{r}$ factors. Thus $H^{\ast}_{\operatorname{\mathrm{tensor}}}(B(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r});\mathbb{Z})=\mathbb{Z}[y_{1},y_{2}]/(ry_{1},ry_{2})$, where $y_{1}=c_{1}(W_{(1,0)})$ and $y_{2}=c_{2}(W_{0,1})$. Next, the Whitney sum formula (see, e.g.,  [21]) shows that $c(V_{1}\oplus V_{2})=c(V_{1})c(V_{2})$ for any two complex $(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r})$-representations $V_{1}$ and $V_{2}$, and one also has $c_{1}(V_{1}\otimes V_{2})=c_{1}(V_{1})+c_{2}(V_{2})$ if both of these representations are $1$-dimensional (see, e.g,  [17]). However, $W_{(i,j)}=W_{(1,0)}^{\otimes i}\otimes W_{(0,1)}^{\otimes j}$ for any $(i,j)$, and as $c(E)=(1+c_{1}(E))$ for any line bundle $E$, the decomposition (7.1) shows that
$c_{N}(W)=[((r-1)/2)!]^{2d}y_{1}^{d(r-1)/2}y_{2}^{d(r-1)/2}\prod_{(a_{1},a_{2})\in A}(a_{1}y_{1}+a_{2}y_{2})^{m},$
where we recall that $A\subset\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}$ consists of all $(a_{1},a_{2})\neq(0,0)$ whose last non-zero coordinate lies in $[(r-1)/2]$.
We have that $c_{N}(W)$ lies in the tensor subring of $H^{\ast}(B(\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}))$, and therefore $c_{N}(E(r,n))=i^{\ast}(c_{N}(W))$ lies in the analogous tensor subring of $H^{\ast}(\overline{X(r,n});\mathbb{Z})$. As $\overline{X_{n}}$ is the $(n-r)$-dimensional skeleton of $B\mathbb{Z}_{r}$, the long exact sequence for the pair $(\overline{X_{n}},B\mathbb{Z}_{r})$ shows that the map $i^{\ast}\colon H^{i}(B\mathbb{Z}_{r})\rightarrow H^{i}(\overline{X_{n}})$ induced from inclusion is an injection if $i\leq n-r$. As $N\leq n-r$, this holds in particular when $i\leq N$. The corresponding map
$\oplus_{i+j\leq N}H^{i}(B\mathbb{Z}_{r};\mathbb{Z})\otimes H^{j}(B\mathbb{Z}_{r};\mathbb{Z})\rightarrow\oplus_{i+j\leq N}H^{i}(\overline{X_{n}};\mathbb{Z})\otimes H^{j}(\overline{X};\mathbb{Z})$
on tensor products is therefore injective as well. Letting $N_{0}=N/2=d\frac{r-1}{2}+m\frac{r^{2}-1}{4}$, it follows that $c_{N}(E(r,n))=i^{*}(c_{N}(W))$ is non-zero provided $c_{N}(W)$ is non-zero when considered as an element of the quotient ring $\mathbb{Z}[ry_{1},ry_{2}]/(ry_{1},ry_{2},y_{1}^{N_{0}+1},y_{2}^{N_{0}+1})$.
Let $A^{\prime}$ be the subset of $A$ consisting of all non-zero elements $(a_{1},a_{2})\in\mathbb{Z}_{r}\oplus\mathbb{Z}_{r}$ whose last non-zero coordinate is equal to 1. Thus $\prod_{(a_{1},a_{2})\in A}(a_{1}y_{1}+a_{2}y_{2})=[(r-1)/2]!\prod_{(a_{1},a_{2})\in A^{\prime}}(a_{1}y_{1}+a_{2}y_{2})^{(r-1)/2}$. On the other hand, $\prod_{(a_{1},a_{2})\in A^{\prime}}(a_{1}y_{1}+a_{2}y_{2})=(y_{1}y_{2}^{r}-y_{2}y_{1}^{r})$ (see, e.g., [34]), so that $c_{n}(W)=cp(y_{1},y_{2})$ where $c$ is a non-zero constant in $\mathbb{Z}_{r}$ and
$p(y_{1},y_{2})=y_{1}^{d(r-1)/2}y_{2}^{d(r-1)/2}(y_{1}y_{2}^{r}-y_{2}y_{1}^{r})^{m(r-1)/2}.$
Now consider the $r$-ary expansion $b_{0}+b_{1}r+\cdots+b_{\ell}r^{\ell}$ of $m^{\prime}:=m(r-1)/2$. By assumption, these coefficients are all even, so $m^{\prime}$ is even as well. Since each of the terms $ky_{1}^{\alpha_{1}}y_{2}^{\alpha_{2}}$ of $p(y_{1},y_{2})$ satisfies $\alpha_{1}+\alpha_{2}=N$, $p(y_{1},y_{2})$ is non-zero when reduced modulo $y_{1}^{N_{0}+1}$ and $y_{2}^{N_{0}+1}$ if and only if the binomial coefficient $\binom{m^{\prime}}{m^{\prime}/2}$ is non-zero in $\mathbb{Z}_{r}$. However, Lucas’s theorem (see, e.g.,  [11]) shows that $\binom{m^{\prime}}{m^{\prime}/2}=\prod_{i=0}^{\ell}\binom{b_{i}}{b_{i}/2}$ when taken in $\mathbb{Z}_{r}$, and therefore $\binom{m^{\prime}}{m^{\prime}/2}$ is non-zero in $\mathbb{Z}_{r}$. Thus $c_{N}(W)$ is non-zero in $\mathbb{Z}[y_{1},y_{2}]/(ry_{1},ry_{2},y_{1}^{N_{0}+1},y_{2}^{N_{0}+1})$, and therefore $c_{N}(E(r,n))\in H^{2N}(\overline{X(r,n)};\mathbb{Z})$ is non-vanishing. This completes the proof of Proposition 6.1.
8. Piercing Distributions for Two Fans
We conclude with a proof of our equidistribution result for two fans. As for a single fan, this follows from a corresponding Dolnikov-type theorem.
Theorem 8.1.
Let $d\geq 1$ be an integer, let $r\geq 3$ be a prime number, and let $m\geq 1$ be an integer such that all of the coefficients in the $r$-ary expansion of $m(r-1)/2$ are even.
  1. (a)
    Suppose that $X$ is a set of $n\geq(r-1)(d+1)+\frac{m(r^{2}-1)}{2}+1$ affinely spanning points in $\mathbb{R}^{n-d-1}$. If $\mathcal{F}$ is a family of subsets of $X$ with $\chi(\mathrm{KG}^{r^{2}}(\mathcal{F}))\leq m$, then there exists a pair of $r$-fans $F_{r}^{1}$ and $F_{r}^{2}$ of codimension $r-2$ whose intersection contains $X$ such that every $F\in\mathcal{F}$ is intersected by two closed half-flats from one of the fans.
  2. (b)
    Suppose that $X$ is a set of $n\geq(r-1)(d+1)+\frac{m(r^{2}-1)}{2}+1$ complex affinely spanning points in $\mathbb{C}^{n-d-1}$. If $\mathcal{F}$ is a family of subsets of $X$ such that $\chi(\mathrm{KG}^{r^{2}}(\mathcal{F}))\leq m$, then there exists a pair of complex regular $r$-fans $F_{r}^{1}$ and $F_{r}^{2}$ in $\mathbb{C}^{n-d-1}$ whose intersection contains $X$ such that every $F\in\mathcal{F}$ is intersected by two closed half-hyperplanes from one of the fans.
Proof of Theorem 1.7.
We prove the real case of Theorem 1.7; the complex case is identical. So let $n\geq(r-1)(d+1)+m\frac{r^{2}-1}{2}$, let $m$ satisfy the condition of Theorem 1.7(a), and let $X=X_{1}\sqcup\cdots\sqcup X_{m}$ be an $m$-coloring of a set $X$ of $n$ affinely spanning points in $\mathbb{R}^{n-d-1}$. Corresponding to the coloring, we let $\mathcal{F}_{k}=\{A\subset X\mid|A\cap X_{k}|>|X_{k}|/r^{2}\}$ for each $k\in[m]$, and we let $\mathcal{F}=\cup_{k=1}^{m}\mathcal{F}_{k}$. We have $\chi(\mathrm{KG}^{r^{2}}(\mathcal{F}))\leq m$ by the pigeon-hole principle, and so by Theorem 8.1 there are two $r$-fans $F_{r}^{1}=\cup_{i=1}^{r}B_{i}^{1}$ and $F_{r}^{2}=\cup_{i=1}^{r}B_{i}^{2}$ of codimension $r-2$ in $\mathbb{R}^{n-d-1}$ such no element of $\mathcal{F}$ is completely contained in any of the intersections $\operatorname{\mathrm{Int}}(B_{1}^{i})\cap\operatorname{\mathrm{Int}}(B_{2}^{j})$. However, $\operatorname{\mathrm{Int}}(B_{1}^{i})\cap\operatorname{\mathrm{Int}}(B_{j}^{2})\cap X$ is a subset of $\operatorname{\mathrm{Int}}(B_{1}^{i})\cap\operatorname{\mathrm{Int}}(B_{2}^{j})$ for any $i,j\in[r]$, so none of the intersections $\operatorname{\mathrm{Int}}(B_{1}^{i})\cap\operatorname{\mathrm{Int}}(B_{2}^{j})$ is a member of $\mathcal{F}$. Thus $|\operatorname{\mathrm{Int}}(B_{1}^{i})\cap\operatorname{\mathrm{Int}}(B_{2}^{j})\cap X_{k}|\leq|X_{k}|/r^{2}$ for all $i,j\in[r]$ and all $k\in[m]$. ∎
Finally, we prove Theorem 8.1. Our proof follows the same procedure as in the case of a single fan, where now we apply Gale duality to the linear case of Theorem 1.8.
Proof of Theorem 8.1.
Let $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$. If $\mathbb{K}=\mathbb{R}$ we let $n\geq(r-1)(d+1)+m\frac{r^{2}-1}{2}+1$, while if $\mathbb{K}=\mathbb{C}$ we let $n\geq(r-1)(2d+1)+m\frac{r^{2}-1}{2}+1$. Suppose now that $X=\{x_{1},\ldots,x_{n}\}$ is a set of $n$ affinely spanning points in $\mathbb{K}^{n-d-1}$ and that $\mathcal{F}$ is family of non-empty subsets of $X$ with $\chi(\mathrm{KG}^{r^{2}}\mathcal{F})\leq m$. As in the proof of Theorem 1.4, we lift $X$ to the hyperplane $\mathbb{K}^{n-d-1}\times\{1\}$ in $\mathbb{K}^{n-d}$ by setting $g_{1}=(x_{1},1),\ldots,g_{n}=(x_{n},1)$. By including the point $g_{n+1}=-\sum_{i=1}^{n}g_{i}$, the sequence $g_{1},\ldots,g_{n+1}$ in $\mathbb{K}^{n-d}$ is the Gale transform of a sequence $a_{1},\ldots,a_{n+1}$ of $n+1$ affinely spanning points in $\mathbb{K}^{d}$. As before, we let $f\colon\Delta_{n}\rightarrow\mathbb{K}^{d}$ be the unique (real) affine linear map which sends each $i\in[n]$ to $a_{i}\in\mathbb{K}^{d}$, and, corresponding to $\mathcal{F}$, we let $\mathcal{F}^{\prime}$ be the family of all subsets $\{i_{1},\ldots,i_{k}\}$ of $[n]$ such that $\{x_{i_{1}},\ldots,x_{i_{k}}\}$ is in $\mathcal{F}$. Thus $\chi(\mathrm{KG}^{r}(\mathcal{F}^{\prime}))\leq m$.
As $n$ is of the correct size, we may apply Theorem 1.8 to the restriction $f\mid_{\Delta_{n-1}}\colon\Delta_{n-1}\rightarrow\mathbb{K}^{d}$ of $f$. This gives two proper Tverberg-tuples $(\sigma_{1},\ldots,\sigma_{r})$ and $(\tau_{1},\ldots,\tau_{r})$ for $f\mid_{\Delta_{n-1}}$ such that none of the sets of the form $\operatorname{\mathrm{Vert}}(\sigma_{i}\cap\tau_{j})$ contains any element of $\mathcal{F}^{\prime}$ as a subset. Now let $I^{1}_{i}=\operatorname{\mathrm{Vert}}(\sigma_{i})$ for all $i\in[r]$ and $I_{j}^{2}=\operatorname{\mathrm{Vert}}(\tau_{j})$ for each $j\in[r]$. By Lemma 3.1 applied to $f$, there exist two linear $r$-fans $F_{r}^{1}=\cup_{i=1}^{r}B_{i}^{1}$ and $F_{r}^{2}=\cup_{j=1}B_{j}^{2}$ in $\mathbb{K}^{n-d}$ such that $g_{1},\ldots,g_{n+1}$ is distributed by both the $I^{1}_{i}$ and the $I^{2}_{j}$. As each $I^{1}_{i}$ and each $I^{j}_{2}$ is a subset of $[n]$, $g_{n+1}$ lies in the center of both fans.
As in the proof of Theorem 1.4, for each $k\in\{1,2\}$ and each $j\in[r]$ we let $(B_{j}^{k})^{\prime}$ be the intersection of $B_{j}^{k}$ with $\mathbb{K}^{n-d-1}\times\{1\}$ and we let $(F_{i}^{k})^{\prime}=\cup_{j=1}^{k}(B_{j}^{k})^{\prime}$. Letting $(B_{j}^{k})^{\prime\prime}$ be the projection of $(B_{j}^{k})^{\prime}$ onto $\mathbb{K}^{n-d-1}\times\{0\}$, the identical argument to the one given there shows that each $(F_{r}^{k})^{\prime\prime}=\cup_{j=1}^{k}(B_{j}^{k})^{\prime\prime}$ is an $r$-fan in $\mathbb{K}^{n-d-1}$ – of codimension $r-2$ when $\mathbb{K}=\mathbb{R}$, and a complex regular one when $\mathbb{K}=\mathbb{C}$. As each $g_{i}=(x_{i},1)$ lies in $(F_{r}^{k})^{\prime}$, we have that $X$ lies in the intersection $(F_{r}^{1})^{\prime\prime}\cap(F_{r}^{2})^{\prime\prime}$ of these fans. Now, for each $i,j\in[r]$ we have that $g_{k}$ lies in $\operatorname{\mathrm{Int}}(B_{i}^{1})\cap\operatorname{\mathrm{Int}}(B_{j}^{2})$ whenever $k\in I^{1}_{j}\cap I^{2}_{j}$. Thus $x_{k}\in\operatorname{\mathrm{Int}}((B_{i}^{1})^{\prime\prime})\cap\operatorname{\mathrm{Int}}((B_{j}^{2})^{\prime\prime})$ for all $k\in I^{1}_{j}\cap I^{2}_{j}$ and all $i,j\in[r]$. We now show that every element of $\mathcal{F}$ is intersected by two half-flats from at least one of the $F_{r}^{k}$, or equivalently that no member of $\mathcal{F}$ is contained in any of the intersections of the form $\operatorname{\mathrm{Int}}((B_{i}^{1})^{\prime\prime})\cap\operatorname{\mathrm{Int}}((B_{j}^{2})^{\prime\prime})$. But if some $F\in F$ were contained in one of the $\operatorname{\mathrm{Int}}((B_{i}^{1})^{\prime\prime})\cap\operatorname{\mathrm{Int}}((B_{j}^{2})^{\prime\prime})$, then $F\times\{1\}$ would be contained in $\operatorname{\mathrm{Int}}(B_{i}^{1})\cap\operatorname{\mathrm{Int}}(B_{j}^{2})$. The $F^{\prime}$ in $\mathcal{F}^{\prime}$ corresponding to $F$ would then be contained in $I_{i}^{1}\cap I_{j}^{2}=\operatorname{\mathrm{Vert}}(\sigma_{i}\cap\tau_{j})$, thereby contradicting Theorem 1.8. Thus none of the intersections $\operatorname{\mathrm{Int}}((B_{i}^{1})^{\prime\prime})\cap\operatorname{\mathrm{Int}}((B_{j}^{2})^{\prime\prime})$ contains any element of $\mathcal{F}$, which concludes the proof.∎
Acknowledgements
The authors are grateful for the support provided by the 2024 Bard Science Research Institute. They also thank Florian Frick for fruitful conversations.